### v1

# D. Leykekhman, B. Vexler and J. Wagner

Abstract
In this paper, we consider a state constrained optimal control problem governed by the transient Stokes equations. The state constraint is given by an $L^{2}$ functional in space, which is required to fulfill a pointwise bound in time. The discretization scheme for the Stokes equations consists of inf-sup stable finite elements in space and a discontinuous Galerkin method in time, for which we have recently established best approximation type error estimates. Using these error estimates, for the discrete control problem we establish error estimates and as a by-product we show an improved regularity for the optimal control. We complement our theoretical analysis with numerical results.
keywords:
keyword1; Keyword2; Keyword3; Keyword4.
1 Introduction
In this paper we consider the following optimal control problem
[…]
subject to the state equation
$\partial_{t}{\bf{u}}-\Delta{\bf{u}}+\nabla p$ $={\bf{q}}$ in $\text{in }I\times\Omega,$ (1.1b)
$\nabla\cdot{\bf{u}}$ $=0$ in $\text{in }I\times\Omega,$
${\bf{u}}$ $={\bf{0}}$ on $\text{on }I\times\partial\Omega,$
${\bf{u}}(0)$ $={\bf{0}}$ in $\text{in }\Omega,$
control constraints
${\bf{q}}_{a}\leq{\bf{q}}(t,x)\leq{\bf{q}}_{b}\quad\text{for almost all }\;(t,x)\in I\times\Omega$ (1.1c)
and state constraints
[…]
Here we assume that $\Omega\subset\mathbb{R}^{d}$, $d=2,3$, is a convex polygonal or polyhedral domain and $I=(0,T]$ is a bounded time interval. The control constraints are given by the constant vectors ${\bf{q}}_{a},{\bf{q}}_{b}\in{(\mathbb{R}\cup\{\pm\infty\})^{d}}$ and satisfy ${\bf{q}}_{a}<{\bf{q}}_{b}$. The constant scalar $\beta$ satisfies $\beta>0$, ${\bf w}(x)$ is a given function in $L^{2}(\Omega)^{d}$ and $\alpha>0$ is some regularization parameter. Note that for ease of presentation, we consider an optimal control problem with homogeneous initial data ${\bf{u}}(0)={\bf{0}}$, while all results also extend to the inhomogeneous case ${\bf{u}}(0)={\bf{u}}_{0}$. The main result of this paper states, that the error between the optimal control $\bar{{\bf{q}}}$ for the continuous problem and the optimal solution $\bar{{\bf{q}}}_{\sigma}$ of the discretized problem satisfies
$\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{\sigma}\rVert_{L^{2}(I;L^{2}(\Omega))}\leq\frac{C}{\sqrt{\alpha}}\,\ln{\frac{T}{k}}(k^{\frac{1}{2}}+h),$
[…]
2 Notation and Preliminary results
We will use the standard notation for the Lebegue and Sobolev spaces over the spatial domain $\Omega$. The pressure space is
[…]
Throughout the paper, vector valued quantities and spaces will be indicated by boldface letters. We denote for a Banach space $X$ and $1\leq p\leq\infty$, by $L^{p}(I;X)$ the Bochner space of $X$-valued functions over $I$, whose $X$-norm is $p$-integrable w.r.t time. If $X$ is reflexive and $1\leq p<\infty$, there holds the following isomorphism, see [20, Corollary 1.3.22]
$L^{p}(I;X)^{*}\cong L^{q}(I;X^{*}),\qquad\text{where }\frac{1}{p}+\frac{1}{q}=1.$ (2.1)
Note that the range includes the value $p=1$ and not $p=\infty$. The dual space of $C(\bar{I};L^{2}(\Omega)^{d})$ is isomorphic to the space of regular $L^{2}(\Omega)^{d}$-valued Borel measures, and we denote it by $\mathcal{M}(\bar{I};L^{2}(\Omega)^{d})\cong(C(\bar{I};{L^{2}(\Omega)^{d}}))^{*}$. Similarly for scalar regular Borel measures of $\bar{I}$, we use the notation $\mathcal{M}(\bar{I})\cong(C(\bar{I}))^{*}$. To denote the vector valued spaces of divergence free functions, having various levels of regularity, we use the following notation
$\mathbf{V}_{0}=\{{\bf{v}}\in L^{2}(\Omega)^{d}:\ \nabla\cdot{\bf{v}}=0\land{\bf{v}}\cdot\mathbf{n}=0\text{ on }\partial\Omega\},\quad\mathbf{V}:=\{{\bf{v}}\in H^{1}_{0}(\Omega)^{d}:\ \nabla\cdot{\bf{v}}=0\},\quad\mathbf{V}_{2}:=H^{2}(\Omega)^{d}\cap\mathbf{V},$
where by ${\bf{v}}\cdot\mathbf{n}$ we denote the generalized normal trace. We denote by $\mathbf{V}^{*}$ the topological dual space of $\mathbf{V}$, and define
$\mathbf{W}$ $:=L^{2}(I;\mathbf{V})\cap H^{1}(I;\mathbf{V}^{*})\hookrightarrow C(\bar{I};{\mathbf{V}_{0}}),$ (2.2)
$\mathbf{X}$ :=$:=\{{\bf{v}}\in L^{2}(I;\mathbf{V}):\partial_{t}{\bf{v}}\in L^{2}(I;\mathbf{V}^{*})+L^{1}(I;\mathbf{V}_{0})\text{ and }{\bf{v}}(0)={\bf{0}}\},$
[…]
We denote by $\left(\cdot,\cdot\right)_{\Omega}$ and $\left(\cdot,\cdot\right)_{I\times\Omega}$ the inner products of $L^{2}(\Omega)^{d}$ and $L^{2}(I;L^{2}(\Omega)^{d})$ respectively, and by $\left\langle\cdot,\cdot\right\rangle_{I\times\Omega}$ the duality pairing between $L^{2}(I;\mathbf{V})$ and $L^{2}(I;\mathbf{V}^{*})$. We introduce the Stokes operator $A$, defined by
$A\colon D(A)\subset\mathbf{V}_{0}\to\mathbf{V}_{0},\ \left(A{\bf{u}},{\bf{v}}\right)_{\Omega}=\left(\nabla{\bf{u}},\nabla{\bf{v}}\right)_{\Omega}.$
The $H^{2}$ regularity results of [9, 23] show that $D(A)=\mathbf{V}_{2}$. As $A$ is a positive, selfadjoint operator, fractional powers $A^{s}$ are well defined. Of special importance is $A^{\frac{1}{2}}$ which is an isometric isomorphism
$A^{\frac{1}{2}}\colon D(A^{\frac{1}{2}})=\mathbf{V}\to\mathbf{V}_{0},$
as it holds $\|A^{\frac{1}{2}}{\bf{u}}\|_{L^{2}(\Omega)}^{2}=\left(A{\bf{u}},{\bf{u}}\right)_{\Omega}=\left(\nabla{\bf{u}},\nabla{\bf{u}}\right)_{\Omega}=\|\nabla{\bf{u}}\|_{L^{2}(\Omega)}^{2}$. For the proof of $D(A^{\frac{1}{2}})=\mathbf{V}$, see [31, Ch. III, Lemma 2.2.1]. By its definition, we can extend $A$ to an operator (denoted by the same symbol) $A\colon\mathbf{V}\to\mathbf{V}^{*}$ yielding another isometric isomporphism between those spaces. Lastly, as
$\|A^{\frac{1}{2}}{\bf{u}}\|_{\mathbf{V}^{*}}=\sup_{{\bf{v}}\in\mathbf{V}}\dfrac{\left\langle A^{\frac{1}{2}}{\bf{u}},{\bf{v}}\right\rangle_{\Omega}}{\|{\bf{v}}\|_{\mathbf{V}}}=\sup_{{\bf{w}}\in\mathbf{V}_{0}}\dfrac{\left\langle A^{\frac{1}{2}}{\bf{u}},A^{-\frac{1}{2}}{\bf{w}}\right\rangle_{\Omega}}{\|A^{-\frac{1}{2}}{\bf{w}}\|_{\mathbf{V}}}=\sup_{{\bf{w}}\in\mathbf{V}_{0}}\dfrac{\left({\bf{u}},{\bf{w}}\right)_{\Omega}}{\|A^{-\frac{1}{2}}{\bf{w}}\|_{\mathbf{V}}}=\sup_{{\bf{w}}\in\mathbf{V}_{0}}\dfrac{\left({\bf{u}},{\bf{w}}\right)_{\Omega}}{\|{\bf{w}}\|_{\mathbf{V}_{0}}}=\|{\bf{u}}\|_{\mathbf{V}_{0}},$
there holds the isometric isomorphism
$A^{\frac{1}{2}}\colon\mathbf{V}_{0}\to\mathbf{V}^{*}.$ (2.3)
Using the above defined function spaces, the weak formulation of the state equation (1.1b) for a given ${\bf{q}}\in L^{1}(I;L^{2}(\Omega)^{d})+L^{2}(I;\mathbf{V}^{*})$ reads as follows: Find ${\bf{u}}\in{\mathbf{X}}$ such that
$\left\langle\partial_{t}{\bf{u}},{\bf{v}}\right\rangle_{\Omega}+\left({\rm NewA}{\bf{u}},{\rm NewA}{\bf{v}}\right)_{\Omega}$ $=\left({\bf{q}},{\bf{v}}\right)_{\Omega}\quad{\text{for all }{\bf{v}}\in\mathbf{V},\ \text{a.e. in }I}.$ (2.4)
For the above weak formulation, there holds the following result.
Theorem 2.1.
For ${\bf{q}}\in L^{1}(I;{L^{2}(\Omega)^{d}})+L^{2}(I;\mathbf{V}^{*})$, there exists a unique solution ${\bf{u}}\in L^{2}(I;\mathbf{V})\cap C(\bar{I};\mathbf{V}_{0})$ solving (2.4) and the following estimate holds
$\|{\bf{u}}\|_{L^{2}(I;\mathbf{V})}+\|{\bf{u}}\|_{L^{\infty}(I;L^{2}(\Omega))}\leq C\|{\bf{q}}\|_{L^{1}(I;\mathbf{V}_{0})+L^{2}(I;\mathbf{V}^{*})}.$
If ${\bf{q}}\in L^{2}(I;{L^{2}(\Omega)^{d}})$, then ${\bf{u}}\in L^{2}(I;{\mathbf{V}_{2}})\cap H^{1}(I;{\mathbf{V}_{0}})\hookrightarrow C(\bar{I};\mathbf{V})$, and there holds
$\|{\bf{u}}\|_{L^{2}(I;H^{2}(\Omega))}+\|{\bf{u}}\|_{L^{\infty}(I;\mathbf{V})}+\|\partial_{t}{\bf{u}}\|_{L^{2}(I;L^{2}(\Omega))}\leq C\|{\bf{q}}\|_{L^{2}(I;L^{2}(\Omega))}.$
Proof.
The first part of this theorem is proven in [32, Chapter III, Theorem 1.1] and the remark on page 179 therein. The $H^{2}$ regularity part can be shown as in [32, Chapter III, Proposition 1.2] using the $H^{2}$ regularity result for the stationary Stokes equations in convex polygonal/polyhedral domains, proven in [9, Theorem 5.5, Theorem 6.3] see also [23, Theorem 2]. ∎
It is a classical result, that the Stokes operator in the Hilbert space setting exhibits maximal parabolic regularity, i.e.
[…]
see [3, Proposition 2.6], holding on fairly general domains, e.g. Lipschitz domains. If $\Omega$ is convex, the $H^{2}$ regularity results further imply that ${\bf{u}}\in L^{p}(I;\mathbf{V}_{2})$. Most often maximal parabolic regularity is treated in $L^{2}$ or $L^{q}$ setting in space, but it can also be extended to settings of weaker spacial regularity. The property (2.3) combined with [2, Lemma 11.4] yields that maximal parabolic regularity also holds in $\mathbf{V}^{*}$, i.e.
[…]
If the right hand side ${\bf{q}}$ is regular enough, there exists an associated pressure to the weak solution of (2.4), formulated in divergence free spaces. The regularity of the pressure depends on the regularity of the right hand side and the velocity component of the solution. There holds the following result, see [3, Theorem 2.10, Corollary 2.11].
Proposition 2.2.
Let ${\bf{q}}\in L^{s}(I;L^{2}(\Omega)^{d})$ for some $1<s<\infty$ and let ${\bf{u}}\in L^{2}(I;\mathbf{V})\cap{C(\bar{I};\mathbf{V}_{0})}$ be the weak solution to (2.4). Then there exists a unique $p\in L^{s}(I;L^{2}_{0}(\Omega))$ such that
$\partial_{t}{\bf{u}}-\Delta{\bf{u}}+\nabla p={\bf{q}},$ (2.7)
which is to be understood as an identity in $L^{s}(I;H^{-1}(\Omega)^{d})$. On convex domains $\Omega$, the pressure satisfies $p\in L^{s}(I;H^{1}(\Omega))$.
3 Continuous Optimal Control Problem
We now introduce the control to state mapping.
Theorem 3.1.
Let $S\colon{{\bf{q}}}\mapsto{{\bf{u}}}$ denote the solution operator for the state equation (2.4). Then $S$ is a bounded linear operator between the following spaces;
  • $S\colon L^{2}(I;\mathbf{V}^{*})\to\mathbf{W}\hookrightarrow C(\bar{I};L^{2}(\Omega)^{d})$
  • $S\colon L^{1}(I;\mathbf{V}_{0})\to C(\bar{I};L^{2}(\Omega)^{d})$
  • $S\colon L^{\infty}(I;L^{2}(\Omega)^{d})\to W^{1,s}(I;L^{2}(\Omega)^{d})\cap L^{s}(I;H^{2}(\Omega)^{d})$,  $1\leq s<\infty$.
Proof.
The first two claims are the direct consequences of Theorem 2.1. Note, that the regularity $\partial_{t}{\bf{u}}\in L^{2}(I;\mathbf{V}^{*})$ can only be obtained by bootstrapping if the right hand side is in $L^{2}(I;\mathbf{V}^{*})$. The third claim is obtained by using the maximal parabolic regularity and $H^{2}$ regularity of the stationary Stokes problem on convex polygonal/polyhedral domains, see [9, Theorem 5.5, Theorem 6.3]. ∎
To abbreviate the notation, we will frequently use ${\bf{u}}({\bf{q}}):=S({\bf{q}})$. Since $S$ is a bounded linear operator between the spaces introduced in the previous theorem, it is Fréchet differentiable, and its directional derivative in direction ${\bf{\delta q}}$ satisfies
$S^{\prime}({\bf{q}})({\bf{\delta q}})=S({\bf{\delta q}}),$
i.e. is independent of ${\bf{q}}$.
Remark 3.2.
The operator $S$ is linear and coincides with its Frechét derivative, due to our choice to work with homogeneous inital data in the state equation (1.1b). The results presented in this work however also hold true in the inhomogeneous initial data case. In this case, the assumption $\beta>0$ then has to be generalized to $\beta>\left({\bf{u}}_{0},{\bf{w}}\right)_{\Omega}$.
Let us next give a characterization of the adjoint operator $S^{*}$. Theorem 2.1, together with [5, Proposition V.1.3], yields that for any ${\bf{q}}\in{\mathbf{Y}}$, there exists a unique solution ${\bf{u}}\in\mathbf{X}$ such that
$a({\bf{u}},{\bf{v}}):=\left\langle\partial_{t}{\bf{u}},{\bf{v}}\right\rangle_{I\times\Omega}+\left(\nabla{\bf{u}},\nabla{\bf{v}}\right)_{I\times\Omega}=\left\langle{\bf{q}},{\bf{v}}\right\rangle_{I\times\Omega}\qquad\forall{\bf{v}}\in{\mathbf{Y}^{*}},$ (3.1)
and the two formulations (2.4) and (3.1) are equivalent. The solution operator $S$ is the inverse of the operator $T\colon\mathbf{X}\to\mathbf{Y}$, $\langle T{\bf{u}},{\bf{v}}\rangle_{\mathbf{Y}\times\mathbf{Y}^{*}}=a({\bf{u}},{\bf{v}})$. As $T$ is invertible, so is $T^{*}$ and its inverse is precisely $S^{*}$. With this construction, the adjoint operator $S^{*}:\mathbf{g}\mapsto{\bf{z}}$, corresponds to the following weak formulation: given $\mathbf{g}\in\mathbf{X}^{*}$, find ${\bf{z}}\in{\mathbf{Y}^{*}}$ satisfying
$\left\langle\partial_{t}{\bf{v}},{\bf{z}}\right\rangle_{I\times\Omega}+\left(\nabla{\bf{v}},\nabla{\bf{z}}\right)_{I\times\Omega}=\left(\mathbf{g},{\bf{v}}\right)_{I\times\Omega}\quad\forall{\bf{v}}\in\mathbf{X}.$ (3.2)
We now discuss the regularity of the adjoint operator.
Corollary 3.3.
Let $S^{*}$ denote the adjoint operator to the solution operator introduced in Theorem 3.1. Then it satisfies
$S^{*}\colon\mathcal{M}(\bar{I};L^{2}(\Omega)^{d})\to L^{2}(I;\mathbf{V})\cap L^{\infty}(I;\mathbf{V}_{0}).$
Proof.
This is a direct consequence of the definition of the adjoint operator, Theorem 3.1, and the isomorphism $(L^{1}(I;\mathbf{V}_{0}))^{*}\cong L^{\infty}(I;\mathbf{V}_{0})$, due to (2.1). ∎
Due to the linearity of the adjoint operator, it again holds
$S^{*}({\bf{\delta q}})=(S^{\prime}({\bf{q}}))^{*}({\bf{\delta q}}).$
For convenience, for the weight ${\bf w}\in L^{2}(\Omega)^{d}$, we define the functional $G_{\bf w}\colon L^{2}(\Omega)^{d}\to\mathbb{R}$ by
$G_{\bf w}({\bf{v}}):=({\bf{v}},{\bf w})_{\Omega}.$ (3.3)
For time-dependent functions ${\bf{v}}\colon I\to L^{2}(\Omega)^{d}$, the application of $G_{{\bf w}}$ is defined by
$G_{\bf w}({\bf{v}})(t):=G_{\bf w}({\bf{v}}(t)).$
Using the functional $G_{\bf w}$, the state constraint (1.1d) can be compactly rewritten as
$G_{\bf w}({\bf{u}})\leq\beta\quad\text{in}\ \bar{I}.$ (3.4)
Remark 3.4.
Due to the continuous embedding $\mathbf{W}\hookrightarrow C(\bar{I};L^{2}(\Omega)^{d})$, we have $G_{\bf w}({\bf{v}})(\cdot)\in C(\bar{I})$ for any ${\bf{v}}\in\mathbf{W}$. Thus we can understand $G_{\bf w}$ as a linear, continuous operator from $\mathbf{W}$ to $C(\bar{I})$.
To write the optimal control problem (1.1a)-(1.1d) into reduced form, we define $\mathcal{G}:=G_{\bf w}\circ S$ and the closed convex cone $\mathcal{K}\subset C(\bar{I})$ by
$\mathcal{K}:=\{v\in C(\bar{I}):\ v\leq\beta\ \text{in}\ \bar{I}\}.$
Using the above definitions, the reduced form reads
$\text{Minimize }\ j({\bf{q}}):=J({\bf{q}},S({\bf{q}}))\quad\text{for}\quad{\bf{q}}\in\mathbf{Q}_{ad}\quad\text{subject to}\quad\mathcal{G}({\bf{q}})\in\mathcal{K},$ (3.5)
where the admissible set is given by
$\mathbf{Q}_{ad}=\{{\bf{q}}\in L^{2}(I;L^{2}(\Omega)^{d}):\quad{\bf{q}}_{a}\leq{\bf{q}}(t,x)\leq{\bf{q}}_{b}\quad\text{a.e. in }I\times\Omega\}.$
We define the projection operator onto the feasible set by
$P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}({\bf{q}}):=\min\{{\bf{q}}_{b},\max\{{\bf{q}},{\bf{q}}_{a}\}\},$
which we understand componentwise and pointwise for every $(t,x)\in I\times\Omega$. Throughout the paper, we will work under the following assumption.
Assumption 1 (Slater condition).
There exists $\tilde{{\bf{q}}}\in\mathbf{Q}_{ad}$ such that $G_{\bf w}({\bf{u}}(\tilde{{\bf{q}}}))<\beta$ for all $t\in\bar{I}$, where ${\bf{u}}(\tilde{{\bf{q}}})$ is the solution of the weak transient Stokes problem (2.4) for this particular control $\tilde{{\bf{q}}}$.
Remark 3.5.
As the homogeneous initial data in our setting necessitate the choice $\beta>0$, if the control constraints ${\bf{q}}_{a},{\bf{q}}_{b}$ admit the control $\tilde{\bf{q}}\equiv{\bf{0}}$, the existence of such a Slater point is immediately given. In that case, the unique solution to the state equation with right hand side $\tilde{\bf{q}}\equiv{\bf{0}}$ is ${\bf{u}}(\tilde{\bf{q}})\equiv{\bf{0}}$ which trivially satisfies $G_{\bf{w}}({\bf{u}}(\tilde{\bf{q}}))=0<\beta$.
Theorem 3.6.
Under 1, there exists a unique optimal control $\bar{\bf{q}}\in L^{2}(I;L^{2}(\Omega)^{d})$ with unique associated state $\bar{\bf{u}}$, solving the optimal control problem (1.1a)-(1.1d).
Proof.
The 1 yields the existence of a feasible $\tilde{\bf{q}}$, such that the associated state $\tilde{\bf{u}}={\bf{u}}(\tilde{\bf{q}})$ satisfies the state constraint. Let us define
[…]
and let $\{{\bf{q}}_{n}\}$ and $\{{\bf{u}}_{n}:={\bf{u}}({\bf{q}}_{n})\}$ denote sequences of feasible controls with associated states, such that $J({\bf{q}}_{n},{\bf{u}}_{n})\to\mathcal{J}$ as $n\to\infty$. As it holds $J({\bf{q}}_{n},{\bf{u}}_{n})\leq{\mathcal{J}+1}$ for large enough $n$, there holds a bound $\|{\bf{q}}_{n}\|_{L^{2}(I;L^{2}(\Omega))}\leq C$ for all $n$. From Theorem 3.1, we obtain $\|{\bf{u}}_{n}\|_{L^{2}(I;H^{2}(\Omega))}+\|{\bf{u}}_{n}\|_{H^{1}(I;L^{2}(\Omega))}\leq C$. We can thus take a subsequence, denoted by the same index, such that
${\bf{q}}_{n}\rightharpoonup\hat{\bf{q}}\text{ in }L^{2}(I;L^{2}(\Omega)^{d}),\quad{\bf{u}}_{n}\rightharpoonup\hat{\bf{u}}\text{ in }L^{2}(I;H^{2}(\Omega)^{d}),\quad\partial_{t}{\bf{u}}_{n}\rightharpoonup\partial_{t}\hat{\bf{u}}\text{in }L^{2}(I;L^{2}(\Omega)^{d}).$
These allow us to pass to the limit in the weak form of the state equation, showing that $\hat{\bf{u}}={\bf{u}}(\hat{\bf{q}})$. Furthermore, as $L^{2}(I;H^{2}(\Omega)^{d})\cap H^{1}(I;L^{2}(\Omega)^{d})\hookrightarrow C({\bar{I}};H^{s}(\Omega)^{d})$ compactly, for $s<1$, see [30, Corollary 8], by taking another subsequence, we obtain ${\bf{u}}_{n}\to\hat{\bf{u}}$ in $C({\bar{I}};L^{2}(\Omega)^{d})$. As $G_{\bf{w}}({\bf{u}}_{n})\leq\beta$ for all $t\in I$, this shows $G_{\bf{w}}(\hat{\bf{u}})\leq\beta$ for all $t\in I$. Lastly, it holds due to the lower semicontinuity of the norms
$J(\hat{\bf{q}},\hat{\bf{u}})\leq\liminf_{n\to\infty}J({\bf{q}}_{n},{\bf{u}}_{n})=\mathcal{J},$
which shows that $\bar{\bf{q}}=\hat{\bf{q}}$ is indeed a minimizer with associated state $\bar{\bf{u}}=\hat{\bf{u}}$. Using uniform convexity of the squared $L^{2}$ norms and linearity of the state equation gives the uniqueness. ∎
Theorem 3.7 (First order optimality system).
Let 1 be fulfilled. Then a control $\bar{{\bf{q}}}\in\mathbf{Q}_{ad}$ with associated state $\bar{{\bf{u}}}={\bf{u}}(\bar{{\bf{q}}})$ is the optimal solution to the problem (1.1a)-(1.1d) if and only if and there exists an adjoint state $\bar{{\bf{z}}}\in L^{2}(I;\mathbf{V})\cap{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{\infty}(I;\mathbf{V}_{0})}$ and a Lagrange multiplier $\bar{\mu}\in(C(\bar{I}))^{*}$ that satisfy:
State equation
$\partial_{t}\bar{{\bf{u}}}-\Delta\bar{{\bf{u}}}+\nabla\bar{p}$ $=\bar{{\bf{q}}}$ in $\text{in }I\times\Omega,$ (3.6a)
$\nabla\cdot\bar{{\bf{u}}}$ $=0$ in $\text{in }I\times\Omega,$
$\bar{{\bf{u}}}$ $={\bf{0}}$ on $\text{on }I\times\partial\Omega,$
$\bar{{\bf{u}}}(0)$ $={\bf{0}}$ in $\text{in }\Omega;$
State constraint and complementarity conditions
$G_{\bf{w}}(\bar{{\bf{u}}})\leq\beta,\quad\bar{\mu}\geq 0\quad\text{and}\quad\langle\bar{\mu},\beta-G_{\bf{w}}(\bar{\bf{u}})\rangle=0;$ (3.6b)
Adjoint equation
$-\partial_{t}\bar{{\bf{z}}}-\Delta\bar{{\bf{z}}}+\nabla\bar{r}$ $=\bar{{\bf{u}}}-{\bf{u}}_{d}+{\bar{\mu}}{\bf{w}}$ in $\text{in }I\times\Omega,$ (3.6c)
$\nabla\cdot\bar{{\bf{z}}}$ $=0$ in $\text{in }I\times\Omega,$
$\bar{{\bf{z}}}$ $={\bf{0}}$ on $\text{on }I\times\partial\Omega,$
$\bar{{\bf{z}}}(T)$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bar{\mu}(\{T\}){\bf{w}}}$ in $\text{in }\Omega;$
Variational inequality
$\left(\alpha\bar{{\bf{q}}}+\bar{{\bf{z}}},\delta{\bf{q}}-\bar{{\bf{q}}}\right)\geq 0\quad\forall\delta{\bf{q}}\in\mathbf{Q}_{ad}\quad\Leftrightarrow\quad\bar{{\bf{q}}}=P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}\left(-\frac{1}{\alpha}\bar{{\bf{z}}}\right).$ (3.6d)
The state (3.6a) and adjoint (3.6c) equations should be understood in the weak sense (2.4) and (3.2) respectively.
Proof.
We derive the optimality conditions from the reduced problem (3.5). By the generalized KKT condition (cf. [6, Theorem 5.2]) and the Slater condition from Assumption 1, i.e., the existence of $\tilde{{\bf{q}}}\in\mathbf{Q}_{ad}$ such that $\mathcal{G}(\tilde{{\bf{q}}})\in\operatorname{int}\mathcal{K}$, the optimality of $\bar{{\bf{q}}}$ is equivalent to the existence of a Lagrange multiplier $\bar{\mu}\in C(\bar{I})^{*}$ and the adjoint state
$\bar{{\bf{z}}}=S^{\prime}(\bar{{\bf{q}}})^{*}(S(\bar{{\bf{q}}})-{\bf{u}}_{d})+\mathcal{G}^{\prime}(\bar{{\bf{q}}})^{*}\bar{\mu}$ (3.7)
satisfying
$\left(\alpha\bar{{\bf{q}}}+\bar{{\bf{z}}},\delta{\bf{q}}-\bar{{\bf{q}}}\right)\geq 0\quad\forall\delta{\bf{q}}\in\mathbf{Q}_{ad}\quad\text{and}\quad\langle v-\mathcal{G}(\bar{{\bf{q}}}),\bar{\mu}\rangle\leq 0\quad\forall v\in\mathcal{K}.$
By definition and linearity of the involved operators, we can write $\mathcal{G}^{\prime}(\bar{\bf{q}})^{*}=S^{*}\circ G_{\mathbf{w}}^{*}$, where $G_{\mathbf{w}}^{*}\colon C(\bar{I})^{*}\to C(\bar{I};L^{2}(\Omega)^{d})^{*}$, $G_{\mathbf{w}}^{*}(\mu)=\mu{\bf{w}}$. Thus it holds $\bar{\bf{z}}=S^{*}(\bar{\bf{u}}-{\bf{u}}_{d}+\bar{\mu}{\bf{w}})$, which gives the proposed regularity of $\bar{\bf{z}}$ as a consequence of Corollary 3.3. To complete the proof, we point to the following equivalence
$(S(\bar{\bf{q}})\in\mathcal{K})\land(\langle v-\mathcal{G}(\bar{{\bf{q}}}),\bar{\mu}\rangle\leq 0\quad\forall v\in\mathcal{K})\quad\Leftrightarrow\quad(G_{{\bf{w}}}(\bar{\bf{u}})\leq\beta)\land(\bar{\mu}\geq 0)\land(\langle\beta-{G}(\bar{{\bf{u}}}),\bar{\mu}\rangle=0),$
which is exactly (3.6b). ∎
From the first order optimality system (3.6a)-(3.6b) we can derive the following regularity results.
Theorem 3.8.
Let $\bar{{\bf{q}}}$ denote the optimal control for the problem (1.1a). Then the following regularity holds:
$\bar{{\bf{q}}}\in L^{2}(I;H^{1}(\Omega)^{d})\cap{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{\infty}(I;L^{2}(\Omega)^{d})}.$
If additionally, ${\bf{q}}_{a},{\bf{q}}_{b}\in\mathbb{R}^{d}$, i.e. are finite, it holds $\bar{{\bf{q}}}\in{L^{\infty}(I\times\Omega)^{d}}$.
Proof.
Since ${\bf w}\in L^{2}(\Omega)^{d}$ and $\bar{\mu}\in(C(\bar{I}))^{*}$, we have that $\bar{\mu}{\bf{w}}\in(C(\bar{I};L^{2}(\Omega)^{d}))^{*}$. By Corollary 3.3, it holds $\bar{\bf{z}}=S^{*}(\bar{\bf{u}}-{\bf{u}}_{d}+\bar{\mu}{\bf{w}})\in L^{2}(I;\mathbf{V})\cap L^{\infty}(I;L^{2}(\Omega)^{d})$. Using that
$\bar{{\bf{q}}}=P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}\left(-\frac{1}{\alpha}\bar{{\bf{z}}}\right),$
and ${\bf{q}}_{a},{\bf{q}}_{b}$ are constant, we have the theorem. ∎
The above result shows the regularity available for the optimal control $\bar{\bf{q}}$ without any additional assumptions. If one assumes higher regularity of the data, especially of the weight ${\bf{w}}$ of the state constraint, we can show improved regularity, and in some cases even a structural result, that the Lagrange multiplier $\bar{\mu}$ contains no Dirac contributions. We first treat a general adjoint problem with measure valued right hand side in Theorem 3.9, before showing regularity of the optimal variables in Theorem 3.10 and the corollaries thereafter.
Theorem 3.9.
Let $\mu\in C(\bar{I})^{*}$ and ${\bf{w}}\in L^{2}(\Omega)^{d}$ be given, and let ${\bf{z}}\in L^{2}(I;\mathbf{V})\cap L^{\infty}(I;\mathbf{V}_{0})$ be the weak solution to
$\left\langle\partial_{t}{\bf{v}},{\bf{z}}\right\rangle_{I\times\Omega}+\left(\nabla{\bf{v}},\nabla{\bf{z}}\right)_{I\times\Omega}=\left\langle{\bf{v}},{\bf{w}}\mu\right\rangle_{I\times\Omega}\quad\text{for all }{\bf{v}}\in\mathbf{X}.$ (3.8)
Then, if additionally ${\bf{w}}\in\mathbf{V}$, it holds ${\bf{z}}\in L^{s}(I;\mathbf{V})\cap\mathrm{BV}(I;\mathbf{V}^{*})$ for all $1\leq s<\infty$. Moreover, if ${\bf{w}}\in\mathbf{V}_{2}$, it holds ${\bf{z}}\in L^{s}(I;\mathbf{V}_{2})\cap\mathrm{BV}(I;\mathbf{V}_{0})$ for all $1\leq s<\infty$.
Proof.
Throughout this proof, let ${\bf{w}}$ satisfy at least ${\bf{w}}\in\mathbf{V}$. Using [1, Theorem 4.31], there exists a normalized function of bounded variation $\tilde{\mu}\in\mathrm{NBV}(\bar{I})$, such that the application of $\mu\in C(\bar{I})^{*}$ to any $\xi\in C(\bar{I})$ can be expressed as a Riemann-Stieltjes integral
$\left\langle\xi,\mu\right\rangle_{I\times\Omega}=\int_{I}\xi(t)\ \mathrm{d}\tilde{\mu}(t).$
Slightly modifying [1, Definition 1.2], we can normalize $\tilde{\mu}$ such that it is of bounded variation, right continuous and satisfies $\tilde{\mu}(T)=0$. If $\xi\in C^{1}(\bar{I})$, due to [1, Proposition 4.24, Theorem 4.17], there holds
$\left\langle\xi,\mu\right\rangle_{I\times\Omega}=-\int_{I}\tilde{\mu}(t)\ \mathrm{d}\xi(t)+\tilde{\mu}(T)\xi(T)-\tilde{\mu}(0)\xi(0)=-\int_{I}\tilde{\mu}(t)\partial_{t}\xi(t)\ \mathrm{d}t-\tilde{\mu}(0)\xi(0)$
where the last integral can be understood equivalently in the Riemann or Lebesgue sense. Working with the Lebesque integral allows us to pass to the limit $C^{1}(\bar{I})\ni\xi_{n}\to\xi\in W^{1,1}(I)\hookrightarrow C(\bar{I})$, showing that
$\left\langle\xi,\mu\right\rangle_{I}=-\left(\tilde{\mu},\partial_{t}\xi\right)_{I}-\tilde{\mu}(0)\xi(0)\qquad\forall\xi\in W^{1,1}(I).$
As due to the definition of $\mathbf{X}$, it holds $\partial_{t}{\bf{v}}\in L^{1}(I;\mathbf{V}_{0})+L^{2}(I;\mathbf{V}^{*})$ and ${\bf{v}}(0)={\bf{0}}$ for all ${\bf{v}}\in\mathbf{X}$, and since ${\bf{w}}\in\mathbf{V}$, we have $\left\langle{\bf{v}},{\bf{w}}\right\rangle_{\Omega}\in W^{1,1}(I)$. Thus we obtain
$\left\langle{\bf{v}},{\bf{w}}\mu\right\rangle_{I\times\Omega}=-\left\langle\tilde{\mu}(t){\bf{w}},\partial_{t}{\bf{v}}\right\rangle_{I\times\Omega}.$ (3.9)
Let us define $\hat{\bf{z}}={\bf{z}}+\tilde{\mu}{\bf{w}}$. Then adding (3.8), (3.9) and the identity $\left(\nabla{\bf{v}},\tilde{\mu}\nabla{\bf{w}}\right)_{I\times\Omega}=\left\langle\tilde{\mu}A{\bf{w}},{\bf{v}}\right\rangle_{I\times\Omega}$ yields
$\left\langle\partial_{t}{\bf{v}},\hat{\bf{z}}\right\rangle_{I\times\Omega}+\left(\nabla{\bf{v}},\nabla\hat{\bf{z}}\right)_{I\times\Omega}=\left\langle\tilde{\mu}A{\bf{w}},{\bf{v}}\right\rangle_{I\times\Omega}\quad\forall{\bf{v}}\in\mathbf{X}.$
As $\tilde{\mu}\in L^{\infty}(I)$ and $\tilde{\mu}(T)<\infty$, $\hat{\bf{z}}$ satisfies a backwards in time Stokes equation with right hand side $\tilde{\mu}A{\bf{w}}$. Depending on the regularity of ${\bf{w}}$, we obtain the following:
Case 1: ${\bf{w}}\in\mathbf{V}$. Here $\tilde{\mu}{\bf{w}}\in\mathrm{BV}(I;\mathbf{V})\hookrightarrow L^{\infty}(I;\mathbf{V})$ and thus $\tilde{\mu}A{\bf{w}}\in L^{\infty}(I;\mathbf{V}^{*})$. Using the maximal parabolic regularity in $\mathbf{V}^{*}$ (2.6), this yields
$\hat{\bf{z}}\in W^{1,s}(I;\mathbf{V}^{*})\cap L^{s}(I;\mathbf{V})\quad\text{for any }1\leq s<\infty.$
Considering the special cases $s=1$ and $s=2$, we obtain
$\hat{\bf{z}}\in W^{1,1}(I;\mathbf{V}^{*})\hookrightarrow\mathrm{BV}(I;\mathbf{V}^{*})\quad\text{and}\quad\hat{\bf{z}}\in H^{1}(I;\mathbf{V}^{*})\cap L^{2}(I;\mathbf{V})\hookrightarrow C(\bar{I};\mathbf{V}_{0}).$
As a consequence ${\bf{z}}=\hat{\bf{z}}-\tilde{\mu}{\bf{w}}\in L^{s}(I;\mathbf{V})\cap\mathrm{BV}(I;\mathbf{V}^{*})$ for any $1\leq s<\infty$.
Case 2: ${\bf{w}}\in\mathbf{V}_{2}$. Here $\tilde{\mu}{\bf{w}}\in\mathrm{BV}(I;\mathbf{V}_{2})\hookrightarrow L^{\infty}(I;\mathbf{V}_{2})$ and thus $\tilde{\mu}A{\bf{w}}\in L^{\infty}(I;\mathbf{V}_{0})$. Using the maximal parabolic regularity in $\mathbf{V}_{0}$ (2.5), this yields
$\hat{\bf{z}}\in W^{1,s}(I;\mathbf{V}_{0})\cap L^{s}(I;\mathbf{V}_{2})\quad\text{for any }1\leq s<\infty.$
Considering the special cases $s=1$ and $s=2$, we obtain
$\hat{\bf{z}}\in W^{1,1}(I;\mathbf{V}_{0})\hookrightarrow\mathrm{BV}(I;\mathbf{V}_{0})\quad\text{and}\quad\hat{\bf{z}}\in H^{1}(I;\mathbf{V}_{0})\cap L^{2}(I;\mathbf{V}_{2})\hookrightarrow C(\bar{I};\mathbf{V}).$
As a consequence ${\bf{z}}=\hat{\bf{z}}-\tilde{\mu}{\bf{w}}\in L^{s}(I;\mathbf{V}_{2})\cap\mathrm{BV}(I;\mathbf{V}_{0})$ for any $1\leq s<\infty$. ∎
Since in the above proof we only have shown continuity of $\hat{\bf{z}}$, but $\tilde{\mu}$ in general can have jumps as it is a $\mathrm{BV}$ function, without further information, we cannot deduce continuity of the whole solution ${\bf{z}}$. For the adjoint equation (3.6c) this translates to the question of continuity in time of $\bar{\bf{z}}$ and by (3.6d) also of $\bar{\bf{q}}$. If in addition to the arguments of the previous theorem, we exploit the information from the optimality system, it turns out, that we can at least show continuity in time of $\bar{\bf{q}}$. As the adjoint state in the optimality system has another term ${\bf{u}}-{\bf{u}}_{d}$ on the right hand side, depending on the available regularity of ${\bf{u}}_{d}$ in time, we might lose some of the time regularity that we just derived.
Theorem 3.10.
Let $(\bar{\bf{q}},\bar{\bf{u}},\bar{\bf{z}},\bar{\mu})$ satisfy the first order necessary optimality conditions (3.6a)-(3.6d). Let additionally ${\bf{u}}_{d}\in L^{s}(I;L^{2}(\Omega)^{d})$ for some $s\in[2,\infty)$ and ${\bf{w}}\in\mathbf{V}$. Then $\bar{\bf{z}}\in L^{s}(I;\mathbf{V})\cap\mathrm{BV}(I;\mathbf{V}^{*})$ and $\bar{\bf{q}}\in L^{s}(I;H^{1}(\Omega)^{d})\cap C(\bar{I};L^{2}(\Omega)^{d})$. If further ${\bf{w}}\in\mathbf{V}_{2}$, then $\bar{\bf{z}}\in L^{s}(I;\mathbf{V}_{2})\cap\mathrm{BV}(I;\mathbf{V}_{0})$ and $\bar{\bf{q}}\in L^{s}(I;W^{1,\infty}(\Omega)^{d})\cap C(\bar{I};H^{1}(\Omega)^{d})$.
Proof.
As in the proof of the previous theorem, we introduce the $\mathrm{NBV}$ function $\tilde{\mu}$ satisfying $\left\langle\xi,\bar{\mu}\right\rangle_{I}=-\left(\tilde{\mu},\partial_{t}\xi\right)_{I}-\tilde{\mu}(0)\xi(0)$ for all $\xi\in W^{1,1}(I)$, and introduce $\hat{\bf{z}}=\bar{\bf{z}}+\tilde{\mu}{\bf{w}}$. This now satisfies
$\left\langle\partial_{t}{\bf{v}},\hat{\bf{z}}\right\rangle_{I\times\Omega}+\left(\nabla{\bf{v}},\nabla\hat{\bf{z}}\right)_{I\times\Omega}=\left({\bf{u}}-{\bf{u}}_{d},{\bf{v}}\right)_{I\times\Omega}+\left\langle\tilde{\mu}A{\bf{w}},{\bf{v}}\right\rangle_{I\times\Omega}\quad\forall{\bf{v}}\in\mathbf{X}.$
The regularity $\bar{\bf{u}}\in L^{\infty}(I;\mathbf{V}_{0})$, the assumption ${\bf{u}}_{d}\in L^{s}(I;L^{2}(\Omega)^{d})$ and the previous arguments again yield that
${\bf{w}}\in\mathbf{V}$ $\Rightarrow\hat{\bf{z}}\in W^{1,s}(I;\mathbf{V}^{*})\cap L^{s}(I;\mathbf{V})\hookrightarrow C(\bar{I};\mathbf{V}_{0})\cap\mathrm{BV}(I;\mathbf{V}^{*})\ \text{ and }\ \bar{\bf{z}}\in L^{s}(I;\mathbf{V}),$
${\bf{w}}\in\mathbf{V}_{2}$ $\Rightarrow\hat{\bf{z}}\in W^{1,s}(I;\mathbf{V}_{0})\cap L^{s}(I;\mathbf{V}_{2})\hookrightarrow C(\bar{I};\mathbf{V})\cap\mathrm{BV}(I;\mathbf{V}_{0})\ \text{ and }\ \bar{\bf{z}}\in L^{s}(I;\mathbf{V}_{2}).$
As in Theorem 3.9 this shows the $\mathrm{BV}$ regularity in time of $\bar{\bf{z}}$. To show the regularity of $\bar{\bf{q}}$, we shall make use of the optimality conditions. The optimal control satisfies
$\bar{\bf{q}}=P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}({\bf{z}}_{\alpha})\quad\text{where}\quad{\bf{z}}_{\alpha}:=-\frac{1}{\alpha}\bar{\bf{z}}=-\frac{1}{\alpha}\hat{\bf{z}}+\frac{1}{\alpha}\tilde{\mu}{\bf{w}}\in C({\bar{I}};\mathbf{V}_{0})+\mathrm{BV}(I;\mathbf{V}).$ (3.10)
The available regularity of $\bar{\bf{z}}$ immediately gives the claimed $L^{s}$ regularity in time of $\bar{\bf{q}}$. Note that due to the application of $P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}$, $\bar{\bf{q}}$ is in general not divergence free anymore and exhibits at most $W^{1,\infty}(\Omega)$ regularity in space, even if $\bar{\bf{z}}$ is smoother. We now turn towards showing continuity in time of $\bar{\bf{q}}$. Let us denote by []delimited-[][$\cdot$][ ⋅ ] the jump function w.r.t. time, i.e. $[\varphi](t)=\varphi(t+)-\varphi(t-)$. Since functions in $\mathrm{BV}(I;X)$ for any Banach space $X$, possess well-defined onesided limits, see [19, Propositions 2.1 & 2.2], we obtain from (3.10) that ${\bf{z}}_{\alpha}(t\pm)\in\mathbf{V}_{0}$ is well defined for any $t\in I$. Hence, by continuity of $P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}\colon L^{2}(\Omega)^{d}\to L^{2}(\Omega)^{d}$ we also have $[\bar{\bf{q}}]=[P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}({\bf{z}}_{\alpha})]\in L^{2}(\Omega)^{d}$ for all $t\in I$. By distinguishing the different cases, it is straightforward to verify, that for any ${\bf{v}}\in C(\bar{I};L^{2}(\Omega)^{d})+\mathrm{BV}(I;L^{2}(\Omega)^{d})$ it holds
$0\leq\left([P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}({\bf{v}})],[P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}({\bf{v}})]\right)_{\Omega}\leq\left([P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}({\bf{v}})],[{\bf{v}}]\right)_{\Omega}\leq\left([{\bf{v}}],[{\bf{v}}]\right)_{\Omega}\quad\text{for all }t\in I.$
Applying this chain of inequalities to ${\bf{v}}={\bf{z}}_{\alpha}$, using $\bar{\bf{q}}=P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}({\bf{z}}_{\alpha})$ and the continuity in time of $\hat{\bf{z}}$, we obtain
$0\leq\|[\bar{\bf{q}}]\|_{L^{2}(\Omega)}^{2}\leq\left([\bar{\bf{q}}],[{\bf{z}}_{\alpha}]\right)_{\Omega}=\frac{[\tilde{\mu}]}{\alpha}\left([\bar{\bf{q}}],{\bf{w}}\right)_{\Omega}.$ (3.11)
For all $t\in I$ satisfying $[\tilde{\mu}](t)=0$, this immediately shows $[\bar{\bf{q}}](t)={\bf{0}}$. Thus let us assume now, that there exists $t^{*}\in I$ with $[\tilde{\mu}](t^{*})\neq 0$. As $\bar{\mu}\geq 0$, there holds for $t_{1}<t_{2}$ due to [1, Theorem 4.17, Proposition 4.24]
$0\leq\left\langle\chi_{[t_{1},t_{2}]},\bar{\mu}\right\rangle_{I}=\int_{[t_{1},t_{2}]}1\mathrm{d}\tilde{\mu}(t)=-\int_{[t_{1},t_{2}]}\tilde{\mu}\ \mathrm{d}1(t)+\tilde{\mu}(t_{2})-\tilde{\mu}(t_{1})=\tilde{\mu}(t_{2})-\tilde{\mu}(t_{1}),$
i.e. $\tilde{\mu}$ is monotonically increasing and therefore, it holds $[\tilde{\mu}]\geq 0$. Moreover it holds
$\bar{\mu}(\{t^{*}\})$ $=\lim_{\varepsilon\to 0}\bar{\mu}((t^{*}-\varepsilon,t^{*}+\varepsilon))=\lim_{\varepsilon\to 0}\left\langle\chi_{(t^{*}-\varepsilon,t^{*}+\varepsilon)},\bar{\mu}\right\rangle_{I}=\lim_{\varepsilon\to 0}\int_{(t^{*}-\varepsilon,t^{*}+\varepsilon)}1\mathrm{d}\tilde{\mu}$ (3.12)
$=\lim_{\varepsilon\to 0}-\int_{(t^{*}-\varepsilon,t^{*}+\varepsilon)}\tilde{\mu}\ \mathrm{d}1+\tilde{\mu}(t^{*}+\varepsilon)-\tilde{\mu}(t^{*}-\varepsilon)=[\tilde{\mu}](t^{*}).$
Thus if $[\tilde{\mu}](t^{*})\neq 0$, then $[\tilde{\mu}](t^{*})>0$ and $t^{*}\in\operatorname{supp}(\bar{\mu})$, yielding that the state constraint is active in $t^{*}$, i.e. $G_{\bf{w}}(\bar{\bf{u}}(t^{*}))=\beta$. As ${\bf{w}}\in\mathbf{V}$, due to (2.4), it holds
$\left(\bar{\bf{q}},{\bf{w}}\right)_{\Omega}+\left(\nabla\bar{\bf{u}},\nabla{\bf{w}}\right)_{\Omega}=\left(\partial_{t}\bar{\bf{u}},{\bf{w}}\right)_{\Omega}=\partial_{t}G_{\bf{w}}(\bar{\bf{u}})\quad\text{a.e. in }I.$
Since $\bar{\bf{q}}\in C(\bar{I};{L^{2}(\Omega)^{d}})+\mathrm{BV}(I;{H^{1}(\Omega)^{d}})$ and $\nabla\bar{\bf{u}}\in C(\bar{I};\mathbf{V})$, this identity shows that $\partial_{t}G_{\bf{w}}(\bar{\bf{u}})\in L^{2}(I)$ has representant with well defined onesided limits, which we shall denote by the same symbol, and it holds
$[\partial_{t}G_{\bf{w}}(\bar{\bf{u}})](t^{*})=\left([\bar{\bf{q}}](t^{*}),{\bf{w}}\right)_{\Omega}\geq\frac{\alpha}{[\tilde{\mu}](t^{*})}\|[\bar{\bf{q}}](t^{*})\|_{L^{2}(\Omega)}^{2}\geq 0,$ (3.13)
where the last inequality holds due to (3.11) and $[\tilde{\mu}](t^{*})>0$. Using $G_{\bf{w}}({\bf{u}})(t)=\int_{0}^{t}\partial_{t}G_{\bf{w}}({\bf{u}})(s)\ \mathrm{d}s$, it is straightforward to check that
$\lim_{h\to 0+}\dfrac{G_{\bf{w}}(\bar{\bf{u}})(t\pm h)-G_{\bf{w}}({\bf{u}})(t)}{h}=\lim_{h\to 0+}\frac{1}{h}\int_{t}^{t\pm h}\partial_{t}G_{\bf{w}}(\bar{\bf{u}})(s)\ \mathrm{d}s=\pm\partial_{t}G_{\bf{w}}(\bar{\bf{u}})(t\pm).$
i.e. the onesided limits of $\partial_{t}G_{\bf{w}}(\bar{\bf{u}})$ correspond to the directional derivatives of $G_{\bf{w}}(\bar{\bf{u}})$. As $G_{\bf{w}}(\bar{\bf{u}})(t^{*})=\beta$, $t^{*}$ is a local maximum of $G_{\bf{w}}(\bar{\bf{u}})$, yielding
±$\pm\partial_{t}G_{\bf{w}}(\bar{\bf{u}})(t^{*}\pm)\leq 0\Rightarrow[\partial_{t}G_{\bf{w}}(\bar{\bf{u}})](t^{*})\leq 0.$
Combining this with (3.13) yields $\|[\bar{\bf{q}}](t^{*})\|_{L^{2}(\Omega)}^{2}=0$, which shows $\bar{\bf{q}}\in C({\bar{I}};L^{2}(\Omega)^{d})$. If additionally ${\bf{w}}\in\mathbf{V}_{2}$, with the same arguments we can show $\bar{\bf{q}}\in C({\bar{I}};H^{1}(\Omega)^{d})$. ∎
Remark 3.11.
In the proof of the above theorem, we can under some circumstances also show continuity in time of $\bar{\bf{z}}$. In fact, the only obstacle preventing this result, is the application of the projection operator onto the control constraints $P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}(\cdot)$. It is in general possible, that at some point in time $\hat{t}$, $\bar{\bf{z}}$ is discontinuous. In such cases, the spacial support of the jump is contained in the set of points $x\in\Omega$, where $-\frac{1}{\alpha}\bar{\bf{z}}(\hat{t}\pm)$ lie outside the set of admissible controls $[{\bf{q}}_{a},{\bf{q}}_{b}]$. As the jump of $\bar{\bf{z}}$ is a scalar multiple of ${\bf{w}}$, this gives a compatibility condition on $\operatorname{supp}({\bf{w}})$ and the set where the control constraints are active. As the active set of the control constraints cannot be known a priori, such a condition on $\bar{\bf{q}}$ is not straightforward to verify. Instead we present in the next two corollaries two sets of assumptions on the data, where we can obtain improved regularity of $\bar{\bf{z}}$ and $\bar{\mu}$.
Corollary 3.12.
Let $(\bar{\bf{q}},\bar{\bf{u}},\bar{\bf{z}},\bar{\mu})$ satisfy the first order necessary optimality conditions (3.6a)-(3.6d). Let further ${\bf{q}}_{a,i}=-\infty$, ${\bf{q}}_{b,i}=+\infty$ for $i=1,...,d$, and ${\bf{w}}\in\mathbf{V}$. Then $\bar{\bf{z}}\in C({\bar{I}};\mathbf{V}_{0})$ and $\bar{\mu}(\{t\})=0$ for all $t\in I$, i.e. $\bar{\mu}$ does not contain any Dirac contributions. If additionally ${\bf{w}}\in\mathbf{V}_{2}$, then $\bar{\bf{z}}\in C({\bar{I}};\mathbf{V})$.
Proof.
This is a direct consequence of the results of Theorem 3.10, as in this case, the identity $\bar{\bf{q}}=-\frac{1}{\alpha}\bar{\bf{z}}$ holds. ∎
Corollary 3.13.
Let $(\bar{\bf{q}},\bar{\bf{u}},\bar{\bf{z}},\bar{\mu})$ satisfy the first order necessary optimality conditions (3.6a)-(3.6d). Let further ${\bf{q}}_{a}<\bf{0}<{\bf{q}}_{b}$, ${\bf{u}}_{d}\in L^{s}(I;L^{2}(\Omega)^{d})$ for some $s>2$, and ${\bf{w}}\in\mathbf{V}_{2}$ satisfy
$\exists x^{*}\in\partial\Omega\text{ s.th. }\forall\varepsilon>0\text{ it holds }\operatorname{supp}({\bf{w}})\cap B_{\varepsilon}(x^{*})\neq\emptyset.$
Then $\bar{\bf{z}}\in C({\bar{I}};\mathbf{V})$ and $\bar{\mu}(\{t\})=0$ for all $t\in I$, i.e. $\bar{\mu}$ does not contain any Dirac contributions.
Proof.
As ${\bf{u}}_{d}\in L^{s}(I;L^{2}(\Omega)^{d})$ for some $s>2$, as in the beginning of the proof of Theorem 3.10, we obtain
$\hat{\bf{z}}\in W^{1,s}(I;\mathbf{V}_{0})\cap L^{s}(I;\mathbf{V}_{2})\hookrightarrow C({\bar{I}};C(\bar{\Omega})),$
where the last embedding holds due to [30, Corollary 8]. As further ${\bf{w}}\in\mathbf{V}_{2}\hookrightarrow C(\bar{\Omega})$, for any $t^{*}\in I$ it thus holds $\bar{\bf{z}}(t^{*}+),\bar{\bf{z}}(t^{*}-)\in C(\bar{\Omega})$. As $\bar{\bf{z}}(t^{*}+)(x^{*})=\bar{\bf{z}}(t^{*}-)(x^{*})={\bf{0}}$, and ${\bf{q}}_{a}<{\bf{0}}<{\bf{q}}_{b}$, there exists $\delta>0$, such that
$-\frac{1}{\alpha}\bar{\bf{z}}(t^{*}+)(x),-\frac{1}{\alpha}\bar{\bf{z}}(t^{*}-)(x)\in({\bf{q}}_{a},{\bf{q}}_{b})\quad\text{for all }x\in B_{\delta}(x^{*})\cap\bar{\Omega}.$
Thus $[\bar{\bf{q}}](t^{*})|_{B_{\delta}(x^{*})}=[-\frac{1}{\alpha}\bar{\bf{z}}](t^{*})|_{B_{\delta}(x^{*})}$. As $\operatorname{supp}({\bf{w}})\cap B_{\delta}(x^{*})\neq\emptyset$, there exists an open subset $\omega\subset B_{\delta}(x^{*})$, such that ${\bf{w}}(x)\neq 0$ for all $x\in\omega$. In the end, we obtain, using the continuity in time of $\bar{\bf{q}}$ shown in Theorem 3.10:
$0=\|[\bar{\bf{q}}](t^{*})\|_{L^{2}(\Omega)}^{2}\geq\int_{\omega}[\bar{\bf{q}}](t^{*})^{2}\mathrm{d}x=\int_{\omega}\left(\left[-\frac{1}{\alpha}\bar{\bf{z}}\right](t^{*})\right)^{2}\mathrm{d}x=\int_{\omega}\frac{[\tilde{\mu}](t^{*})^{2}}{\alpha^{2}}|{\bf{w}}|^{2}\mathrm{d}x\geq 0$
As $|{\bf{w}}|>0$ for all $x\in\omega$, this shows $[\tilde{\mu}](t^{*})=0$. As $t^{*}$ was arbitrary, this concludes the proof. ∎
4 Finite Element Approximation of the State Equation
4.1 Spatial discretization
Let $\{\mathcal{T}_{h}\}$ be a family of triangulations of $\bar{\Omega}$, consisting of closed simplices, where we denote by $h$ the maximum mesh-size. Let $\mathbf{X}_{h}\subset H^{1}_{0}(\Omega)^{d}$ and $M_{h}\subset L^{2}_{0}(\Omega)$ be a pair of compatible finite element spaces, i.e., there holds a uniform discrete inf-sup condition,
$\sup_{{\bf{v}}_{h}\in\mathbf{X}_{h}}\frac{\left(q_{h},\nabla\cdot{\bf{v}}_{h}\right)_{\Omega}}{\lVert\nabla{\bf{v}}_{h}\rVert_{L^{2}(\Omega)}}\geq\gamma\|{q_{h}}\|_{L^{2}(\Omega)}\quad\forall q_{h}\in M_{h},$ (4.1)
with a constant $\gamma>0$ independent of $h$. We shall work under the assumption that the discrete spaces have the following approximative properties.
Assumption 2.
There exist interpolation operators $i_{h}\colon H^{2}(\Omega)^{d}\cap H^{1}_{0}(\Omega)^{d}\to\mathbf{X}_{h}$ and $r_{h}\colon L^{2}(\Omega)\to M_{h}$, such that
$\|\nabla({\bf{v}}-i_{h}{\bf{v}})\|_{L^{2}(\Omega)}$ $\leq ch\|\nabla^{2}{\bf{v}}\|_{L^{2}(\Omega)}$ $\forall{\bf{v}}\in H^{2}(\Omega)^{2}\cap H^{1}_{0}(\Omega)^{2},$
$\|q-r_{h}q\|_{L^{2}(\Omega)}$ $\leq ch\|\nabla q\|_{L^{2}(\Omega)}$ $\forall q\in H^{1}(\Omega).$
This assumption is valid, for example, for Taylor-Hood and MINI finite elements on shape regular meshes, see [3, Assumption 7.2]. We introduce the vectorial discrete Laplace operator $-\Delta_{h}\colon\mathbf{X}_{h}\rightarrow\mathbf{X}_{h}$ by
$(-\Delta_{h}{\bf{v}}_{h},\chi)=(\nabla{\bf{v}}_{h},\nabla\chi),\quad\forall{\bf{v}}_{h},\chi\in\mathbf{X}_{h}.$ (4.2)
We also define the space of discretely divergence-free vector fields $\mathbf{V}_{h}$ as
$\mathbf{V}_{h}=\Set{{\bf{v}}_{h}\in\mathbf{X}_{h}\colon(\nabla\cdot{\bf{v}}_{h},q_{h})=0\quad\forall q_{h}\in M_{h}}.$ (4.3)
While on a computational level, especially for the examples presented in Section 8, we work with a discrete velocity-pressure formulation, in our theoretical analysis we will always use the equivalent formulation in discretely divergence free spaces, in order to shorten notation. One exception is the following stationary Stokes problem of finding for some given $f\in H^{-1}(\Omega)^{d}$ a solution $({\bf{u}},p)\in H^{1}_{0}(\Omega)^{d}\times L^{2}_{0}(\Omega)$ to
$\left(\nabla{\bf{u}},\nabla{\bf{v}}\right)_{\Omega}-\left(\nabla\cdot{\bf{v}},p\right)_{\Omega}+\left(\nabla\cdot{\bf{u}},q\right)_{\Omega}=\left({\bf{f}},{\bf{v}}\right)_{\Omega}\quad\forall({\bf{v}},q)\in H^{1}_{0}(\Omega)^{d}\times L^{2}_{0}(\Omega).$ (4.4)
Its discrete approximation in velocity-pressure formulation reads: Find $({\bf{u}}_{h},p_{h})\in\mathbf{X}_{h}\times M_{h}$ satisfying
$\left(\nabla{\bf{u}}_{h},\nabla{\bf{v}}_{h}\right)_{\Omega}-\left(\nabla\cdot{\bf{v}}_{h},p_{h}\right)_{\Omega}+\left(\nabla\cdot{\bf{u}}_{h},q_{h}\right)_{\Omega}=\left({\bf{f}},{\bf{v}}_{h}\right)_{\Omega}\quad\forall({\bf{v}}_{h},q_{h})\in\mathbf{X}_{h}\times M_{h}.$ (4.5)
The above discrete system can be interpreted as a Stokes Ritz projection: given $({\bf{u}},p)\in H^{1}_{0}(\Omega)^{d}\times L^{2}_{0}(\Omega)$, find $(R_{h}^{S}({\bf{u}},p),R_{h}^{S,p}({\bf{u}},p)):=({\bf{u}}_{h},p_{h})\in\mathbf{X}_{h}\times M_{h}$, satisfying
$\left(\nabla({\bf{u}}-{\bf{u}}_{h}),\nabla{\bf{v}}_{h}\right)_{\Omega}-\left(\nabla\cdot{\bf{v}}_{h},(p-p_{h})\right)_{\Omega}+\left(\nabla\cdot({\bf{u}}-{\bf{u}}_{h}),q_{h}\right)_{\Omega}=0\quad\forall({\bf{v}}_{h},q_{h})\in\mathbf{X}_{h}\times M_{h}.$ (4.6)
Note that, if ${\bf{u}}\in\mathbf{V}$, then it holds $R_{h}^{S}({\bf{u}},p)\in\mathbf{V}_{h}$. Further, the Stokes Ritz projection satisfies the following stability, see [4, Theorem 5.2.1].
$\|\nabla R_{h}^{S}({\bf{u}},p)\|_{L^{2}(\Omega)}+\|R_{h}^{S,p}({\bf{u}},p)\|_{L^{2}(\Omega)}\leq C(\|\nabla{\bf{u}}\|_{L^{2}(\Omega)}+\|p\|_{L^{2}(\Omega)}).$
Let us recall the following error estimates for the stationary discrete Stokes problem: See [4, Theorem 5.25], [22, Theorems 4.21, 4.25, 4.28], [15, Theorems 53.17 & 53.19] or [17, Chapter II, Theorems 1.8 & 1.9].
Theorem 4.1.
Let $({\bf{u}},p)$ and $({\bf{u}}_{h},p_{h})$ denote the solutions to the continuous and discrete stationary Stokes problems (4.4) and (4.5) respectively. Then there holds the estimate
$\|{\bf{u}}-{\bf{u}}_{h}\|_{H^{1}(\Omega)}+\|p-p_{h}\|_{L^{2}(\Omega)}\leq Ch(\|{\bf{u}}\|_{H^{2}(\Omega)}+\|p\|_{H^{1}(\Omega)}).$
If $\Omega$ is convex, there further holds the estimate
$\|{\bf{u}}-{\bf{u}}_{h}\|_{L^{2}(\Omega)}\leq Ch(\|{\bf{u}}-{\bf{u}}_{h}\|_{H^{1}(\Omega)}+\|p-p_{h}\|_{L^{2}(\Omega)}).$
Note, that while [4, Theorem 5.5.6] contains formal results, on how to derive error estimates also for $p$ in a weaker norm, e.g. $H^{-1}(\Omega)$, the argument requires $H^{2}$ regularity results for the compressible Stokes equations. As the corresponding results in [23, 9] require an additional decaying condition for the compressibility data, this makes derivation of weaker error estimates for the pressure complicated.
4.2 Temporal discretization: the discontinuous Galerkin method
In this section we introduce the discontinuous Galerkin method for the time discretization of the transient Stokes equations, a similar method was considered, e.g., in [8]. For that, we partition $I=(0,T]$ into subintervals $I_{m}=(t_{m-1},t_{m}]$ of length ${k}_{m}=t_{m}-t_{m-1}$, where $0=t_{0}<t_{1}<\dots<t_{M-1}<t_{M}=T$. The maximal and minimal time steps are denoted by ${k}=\max_{m}{k}_{m}$ and ${k}_{\min}=\min_{m}{k}_{m}$, respectively. The time partition fulfills the following assumptions:
  1. 1.
    There are constants $C,{\theta}>0$ independent of ${k}$ such that
    ${k}_{\min}\geq C{k}^{{\theta}}.$ (4.7)
  2. 2.
    There is a constant $\kappa>0$ independent of ${k}$ such that for all
    $m=1,2,\dots,M-1$
    $\kappa^{-1}\leq\frac{{k}_{m}}{{k}_{m+1}}\leq\kappa.$ (4.8)
  3. 3.
    It holds ${k}\leq\frac{T}{4}$.
For a given Banach space $\mathcal{B}$, we define the semi-discrete space $X_{k}^{0}(\mathcal{B})$ of piecewise constant functions in time as
$X_{k}^{0}(\mathcal{B})=\Set{{\bf{v}}_{k}\in L^{2}(I;\mathcal{B})\colon{\bf{v}}_{k}}{{}_{I_{m}}\equiv{\bf{v}}_{m}\text{ for some }{\bf{v}}_{m}\in\mathcal{B},m=1,2,\dots,M}.$ (4.9)
We use the following standard notation for a function ${\bf{u}}\in X_{k}^{r}(\mathcal{B})$ to denote one-sided limits and jumps at the time nodes
${\bf{u}}_{m}^{+}=\lim_{\varepsilon\rightarrow 0^{+}}{\bf{u}}(t_{m}+\varepsilon),\quad{\bf{u}}_{m}^{-}=\lim_{\varepsilon\rightarrow 0^{+}}{\bf{u}}(t_{m}-\varepsilon),\quad[{\bf{u}}]_{m}={\bf{u}}_{m}^{+}-{\bf{u}}_{m}^{-}.$ (4.10)
We define the bilinear form ${B}(\cdot,\cdot)$ by
${B}({\bf{u}},{\bf{v}})={\sum_{m=1}^{M}(\partial_{t}{\bf{u}},{\bf{v}})_{I_{m}\times\Omega}}+(\nabla{\bf{u}},\nabla{\bf{v}})_{I\times\Omega}+\sum_{m=2}^{M}([{\bf{u}}]_{m-1},{\bf{v}}^{+}_{m-1})_{\Omega}+({\bf{u}}_{0}^{+},{\bf{v}}_{0}^{+})_{\Omega}.$
With this bilinear form, we define the fully discrete approximation for the transient Stokes problem on the discretely divergence free space $X_{k}^{0}(\mathbf{V}_{h})$:
${\bf{u}}_{kh}\in X_{k}^{0}(\mathbf{V}_{h})\;:\;{B}({\bf{u}}_{kh},{\bf{v}}_{kh})=({\bf{q}},{\bf{v}}_{kh})_{I\times\Omega}\quad\forall{\bf{v}}_{kh}\in X_{k}^{0}(\mathbf{V}_{h}).$ (4.11)
The unique solution to this system is stable, as the following theorem summarizes:
Theorem 4.2.
Let ${\bf{q}}\in L^{2}(I;H^{-1}(\Omega)^{d})+L^{1}(I;L^{2}(\Omega)^{d})$. Then there exists a unique solution ${\bf{u}}_{kh}\in X^{0}_{k}(\mathbf{V}_{h})$ of (4.11), satsifying
$\|{\bf{u}}_{kh}\|_{L^{2}(I;H^{1}(\Omega))}+\|{\bf{u}}_{kh}\|_{L^{\infty}(I;L^{2}(\Omega))}\leq C\|{\bf{q}}\|_{L^{2}(I;H^{-1}(\Omega))+L^{1}(I;L^{2}(\Omega))}.$
Proof.
As (4.11) poses a square system of linear equations in finite dimensions, it suffices to show uniqueness. This is a standard argument, making use of a discrete Gronwall Lemma, see e.g. [34, Theorem 4.13] for a proof focusing especially on the $L^{1}(I;L^{2}(\Omega)^{d})$ right hand side case. ∎
Remark 4.3.
Rearranging terms in the definition of the bilinear form gives the following dual representation of $B(\cdot,\cdot)$
$B({\bf{u}},{\bf{v}})={-\sum_{m=1}^{M}({\bf{u}},\partial_{t}{\bf{v}})_{I_{m}\times\Omega}}+(\nabla{\bf{u}},\nabla{\bf{v}})_{I\times\Omega}-\sum_{m=1}^{M-1}({\bf{u}}_{m}^{-},[{\bf{v}}]_{m})_{\Omega}+({\bf{u}}_{M}^{-},{\bf{v}}_{M}^{-})_{\Omega}.$
With the same arguments as above, for given $\mathbf{g}\in L^{2}(I;H^{-1}(\Omega)^{d})+L^{1}(I;L^{2}(\Omega)^{d})$, solutions ${\bf{z}}_{kh}\in X^{0}_{k}(\mathbf{V}_{h})$ to the discrete dual equation
$B({\bf{v}}_{kh},{\bf{z}}_{kh})=\left(\mathbf{g},{\bf{v}}_{kh}\right)_{I\times\Omega}\quad\forall{\bf{v}}_{kh}\in X_{k}^{0}(\mathbf{V}_{h})$
exist, are unique and satisfy the stability
$\|{\bf{z}}_{kh}\|_{L^{2}(I;H^{1}(\Omega))}+\|{\bf{z}}_{kh}\|_{L^{\infty}(I;L^{2}(\Omega))}\leq C\|\mathbf{g}\|_{L^{2}(I;H^{-1}(\Omega))+L^{1}(I;L^{2}(\Omega))}.$
4.3 Best approximation type fully discrete error estimate for the Stokes problem in $L^{\infty}(I;L^{2}(\Omega)^{d})$ norm
In our recent paper [3], we have established a best approximation type error estimate for the Stokes problem in the $L^{\infty}(I;L^{2}(\Omega)^{d})$ norm. From this more general result, we obtain in the case of homogeneous initial data the following result, see [3, Corollary 6.4].
Theorem 4.4.
Let ${\bf{q}}\in L^{s}(I;L^{2}(\Omega)^{d})$ for some $s>1$ and let ${\bf{u}}\in\mathbf{W}$ be the weak solution to (2.4) with associated pressure $p$ in the sense of (2.7). Let ${\bf{u}}_{kh}\in X_{k}^{0}(\mathbf{V}_{h})$ be the fully discrete Galerkin solution to (4.11). Then there exists a constant $C$ independent of $k$ and $h$, such that for any $\chi\in X_{k}^{0}(\mathbf{V}_{h})$ there holds
$\|{\bf{u}}-{\bf{u}}_{kh}\|_{L^{\infty}(I;L^{2}(\Omega))}\leq C\ell_{k}\left(\|{\bf{u}}-\chi\|_{L^{\infty}(I;L^{2}(\Omega))}+\|{\bf{u}}-R^{S}_{h}({\bf{u}},p)\|_{L^{\infty}(I;L^{2}(\Omega))}\right),$
where $\ell_{k}=\ln{\frac{T}{k}}$ and $R^{S}_{h}({\bf{u}},p)$ is the stationary finite element Stokes projection introduced in (4.6).
Using the error estimates for the stationary Stokes Ritz projection of Theorem 4.1, in [3, Theorem 7.4] the following estimate in terms of explicit orders of convergence was shown.
Corollary 4.5.
If in addition to assumptions of Theorem 4.4, the domain $\Omega$ is convex, and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{q}}}\in L^{\infty}(I;L^{2}(\Omega)^{d})$, then there exists a constant $C$ independent of $k$ and $h$ such that
$\|{\bf{u}}-{\bf{u}}_{kh}\|_{L^{\infty}(I;L^{2}(\Omega))}\leq C\ell_{k}^{2}(h^{2}+k)\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{q}}}\|_{L^{\infty}(I;L^{2}(\Omega))}.$
The above results are valid in $L^{2}(I;L^{2}(\Omega)^{d})$ norm as well. However, using the energy and duality arguments it is possible to show the corresponding results log-free and with less regularity assumptions on the data (cf. [24, Theorem 11 & Theorem 13]).
Theorem 4.6.
Let ${\bf{q}}\in L^{2}(I;L^{2}(\Omega)^{d})$ and let ${\bf{u}}\in\mathbf{W}$ be the weak solution to (2.4) with associated pressure $p$ in the sense of (2.7). Let ${\bf{u}}_{kh}\in X_{k}^{0}(\mathbf{V}_{h})$ be the fully discrete Galerkin solution to (4.11). Then there exists a constant $C$, independent of $k$ and $h$, such that for any $\chi\in X_{k}^{0}(\mathbf{V}_{h})$, there holds
$\|{\bf{u}}-{\bf{u}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}\leq C\left(\|{\bf{u}}-\chi\|_{L^{2}(I;L^{2}(\Omega))}+\|{\bf{u}}-R^{S}_{h}({\bf{u}},p)\|_{L^{2}(I;L^{2}(\Omega))}+\|{\bf{u}}-\pi_{k}{\bf{u}}\|_{L^{2}(I;L^{2}(\Omega))}\right),$ (4.12a)
and
$\|{\rm NewA}({\bf{u}}-{\bf{u}}_{kh})\|_{L^{2}(I;L^{2}(\Omega))}\leq C\bigg{(}\|{\rm NewA}({\bf{u}}-\chi)\|_{L^{2}(I;L^{2}(\Omega))}$ $+\|{\rm NewA}({\bf{u}}-R^{S}_{h}({\bf{u}},p))\|_{L^{2}(I;L^{2}(\Omega))}$ (4.12b)
$+\|{\rm NewA}({\bf{u}}-\pi_{k}{\bf{u}})\|_{L^{2}(I;L^{2}(\Omega))}\bigg{)},$
where $R^{S}_{h}({\bf{u}},p)$ is the stationary finite element Stokes projection defined in (4.6) and $\pi_{k}$ is the time projection onto $X^{0}_{k}(\mathbf{V})$, with $\pi_{k}v|_{I_{m}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=v(t_{m}^{-})}$ for $m=1,2,\dots,M$.
Corollary 4.7.
If in addition to assumptions of Theorem 4.6, the domain $\Omega$ is convex, then there exists a constant $C$ independent of $k$ and $h$ such that
$\|{\bf{u}}-{\bf{u}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}\leq C(h^{2}+k)\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{q}}}\|_{L^{2}(I;L^{2}(\Omega))}.$
5 Variational Discretization of the Optimal Control Problem
In this section we consider the optimal control problem subject to the fully discretized Stokes equations. We consider a variational discretization for the controls, i.e., do not fix a finite dimensional approximation of the control space yet. The problem reads
Minimize $\text{Minimize }\;J({\bf{q}}_{kh},{\bf{u}}_{kh})=\frac{1}{2}\lVert{\bf{u}}_{kh}-{\bf{u}}_{d}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\frac{\alpha}{2}\lVert{\bf{q}}_{kh}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}$ (5.1a)
over all ${\bf{q}}_{kh}\in\mathbf{Q}_{ad}$, ${\bf{u}}_{kh}\in X^{0}_{k}(\mathbf{V}_{h})$, subject to
$B({\bf{u}}_{kh},{\bf{v}}_{kh})=({\bf{q}}_{kh},{\bf{v}}_{kh})_{I\times\Omega}\quad\text{for all}\quad{\bf{v}}_{kh}\in X^{0}_{k}(\mathbf{V}_{h})$ (5.1b)
$G_{\bf{w}}({\bf{u}}_{kh})\mid_{I_{m}}\leq\beta\quad\text{for}\quad m=1,2,\dots,M.$ (5.1c)
Following the structure of Section 3 and using Theorem 4.2, we introduce the discrete analog to the control state map,
[…]
The finitely many state constraints we describe with the help of the continuous linear operator $\mathcal{G}_{kh}\colon\mathbf{Q}_{ad}\to\mathbb{R}^{M}$ with $(\mathcal{G}_{kh}({\bf{q}}))_{m}:=G_{\bf w}\circ S_{kh}({\bf{q}})\mid_{I_{m}}$ for $m=1,2,\dots,M$. Using the set
$\mathcal{K}_{kh}:=\{{\bf{v}}\in\mathbb{R}^{M}\ \mid\ {\bf{v}}_{m}\leq b,\ m=1,2,\dots,M\},$
we can rewrite the problem (5.1a)-(5.1c) in the reduced form:
$\text{Minimize }\ j_{kh}({\bf{q}}_{kh}):=J({\bf{q}}_{kh},S_{kh}({\bf{q}}_{kh}))\quad\text{over}\quad{\bf{q}}_{kh}\in\mathbf{Q}_{ad}\quad\text{subject to}\quad\mathcal{G}_{kh}({\bf{q}}_{kh})\in\mathcal{K}_{kh}.$ (5.3)
Before discussing wellposedness and optimality conditions of this discrete problem, we shall show, that the Slater assumption on the continuous level carries over to the discrete problem. As we achieve this with the finite element error estimates presented in Section 4, we need to impose a rather weak coupling condition between $k$ and $h$, allowing us to deduce convergence in the result of Corollary 4.5. Throughout the remainder of this work, we thus work under the following assumption.
Assumption 3.
There exists a function $\Phi\colon(0,1)\to(0,\infty)$ with $\lim_{h\to 0}\Phi(h)=0$, such that the discretization parameters $k$ and $h$ satisfy
$\left|\ln\left(\frac{T}{k}\right)\right|h\leq\Phi(h).$
Remark 5.1.
This assumption is valid, e.g. if there exists a constant $C>0$ such that $\left|\ln\left(\frac{T}{k}\right)\right|h|\ln h|\leq C$. As the choice of the term $|\ln h|$ in such a condition can be made arbitrarily weak, we have chosen to work under the more general formulation of 3.
Lemma 5.2.
There exists $h_{0}>0$ such that for any $h\leq h_{0}$ and $k$ satisfying 3, the Slater point $\tilde{\bf{q}}\in\mathbf{Q}_{ad}$ from 1 satisfies the following discrete Slater condition
$G_{\bf{w}}({\bf{u}}_{kh}(\tilde{q}))<\beta{\qquad\text{for all }t\in\bar{I}.}$
Proof.
Using that $G_{\bf w}({\bf{u}}(\tilde{{\bf{q}}}))<\beta$ in $\bar{I}$, by the Slater condition Assumption 1 there exists $\delta>0$ such that $G_{\bf w}({\bf{u}}(\tilde{{\bf{q}}}))\leq\beta-\delta$. For arbitrary $\hat{\bf{q}}\in C^{\infty}(I\times\Omega)^{d}$ it holds due to triangle inequality
$\|{\bf{u}}(\tilde{\bf{q}})-{\bf{u}}_{kh}(\tilde{\bf{q}})\|_{L^{\infty}(I;L^{2}(\Omega))}\leq\|{\bf{u}}(\tilde{\bf{q}})-{\bf{u}}_{kh}(\tilde{\bf{q}})-{\bf{u}}(\hat{\bf{q}})+{\bf{u}}_{kh}(\hat{\bf{q}})\|_{L^{\infty}(I;L^{2}(\Omega))}+\|{\bf{u}}(\hat{\bf{q}})-{\bf{u}}_{kh}(\hat{\bf{q}})\|_{L^{\infty}(I;L^{2}(\Omega))}.$
Using the continuous and fully discrete stability results of the state equations, presented in Theorem 2.1 and Theorem 4.2, as well as the error estimate Corollary 4.5 for the problem with right hand side $\hat{\bf{q}}$, we obtain
$\|{\bf{u}}(\tilde{\bf{q}})-{\bf{u}}_{kh}(\tilde{\bf{q}})\|_{L^{\infty}(I;L^{2}(\Omega))}\leq C\|\tilde{\bf{q}}-\hat{\bf{q}}\|_{L^{2}(I\times\Omega)}+C\ell_{k}^{2}(k+h^{2})\|\hat{\bf{q}}\|_{L^{\infty}(I;L^{2}(\Omega))}.$
For any $\varepsilon>0$, due to the density of $C^{\infty}(I\times\Omega)^{d}$ in ${L^{2}(I;L^{2}(\Omega)^{d})}$, we can find $\hat{\bf{q}}_{\varepsilon}$ such that $C\|\tilde{\bf{q}}-\hat{\bf{q}}_{\varepsilon}\|_{L^{2}(I\times\Omega))}<\frac{\varepsilon}{2}$. Moreover, for $h\leq h_{0}$ sufficiently small, and $k$ satisfying 3, it also holds $C\ell_{k}^{2}(k+h^{2})\|\hat{\bf{q}}_{\varepsilon}\|_{L^{\infty}(I;L^{2}(\Omega))}<\frac{\varepsilon}{2}$. Thus in total $\|{\bf{u}}(\tilde{\bf{q}})-{\bf{u}}_{kh}(\tilde{\bf{q}})\|_{L^{\infty}(I;L^{2}(\Omega))}<\varepsilon$. Choosing $\varepsilon$ small enough, such that $\|{\bf{w}}\|_{L^{2}(\Omega)}\varepsilon<\delta$, we obtain
$G({\bf{u}}_{kh}(\tilde{{\bf{q}}}))=G({\bf{u}}(\tilde{{\bf{q}}}))+G({\bf{u}}_{kh}(\tilde{{\bf{q}}})-{\bf{u}}(\tilde{{\bf{q}}}))\leq\beta-\delta+\|{\bf{w}}\|_{L^{2}(\Omega)}\|{\bf{u}}(\tilde{{\bf{q}}})-{\bf{u}}_{kh}(\tilde{{\bf{q}}})\|_{L^{\infty}(I;L^{2}(\Omega))}<\beta.$ (5.4)
Theorem 5.3.
Let $k$ and $h$ satisfy 3, and let $h$ be small enough. Then there exists a unique solution $(\bar{\bf{q}}_{kh},\bar{\bf{u}}_{kh})$ to the optimal control problem (5.1a)-(5.1c).
Proof.
As Lemma 5.2 shows feasibility of $(\tilde{\bf{q}},{\bf{u}}_{kh}(\tilde{\bf{q}}))$ under the given assumptions, the existence proof follows the same steps as the one of Theorem 3.6 on the continuous level. ∎
Theorem 5.4 (Discrete first order optimality system).
A control $\bar{{\bf{q}}}_{kh}\in\mathbf{Q}_{ad}$ and the associated state $\bar{{\bf{u}}}_{kh}={\bf{u}}_{kh}(\bar{{\bf{q}}}_{kh})\in X^{0}_{k}(\mathbf{V}_{h})$ is the optimal solution to the problem (5.1a)-(5.1c) if and only if there exists an adjoint state $\bar{{\bf{z}}}_{kh}\in X^{0}_{k}(\mathbf{V}_{h})$ and a Lagrange multiplier $\bar{\mu}_{kh}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{1}(I)}$ that satisfy:
Discrete state equation
$B(\bar{{\bf{u}}}_{kh},{\bf{v}}_{kh})=(\bar{{\bf{q}}}_{kh},{\bf{v}}_{kh})_{I\times\Omega},\quad\forall{\bf{v}}_{kh}\in X^{0}_{k}(\mathbf{V}_{h});$ (5.5a)
Discrete state constraint and complementarity conditions
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}G_{\bf{w}}(\bar{\bf{u}}_{kh})|_{I_{m}}\leq\beta,\ m=1,...,M},\quad\bar{\mu}_{kh}\geq 0\quad\text{and}\quad\langle\bar{\mu}_{kh},b-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}G_{\bf{w}}(\bar{\bf{u}}_{kh})}\rangle=0;$ (5.5b)
Discrete adjoint equation
$B({\bf{v}}_{kh},\bar{{\bf{z}}}_{kh})=(\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d}+{\bar{\mu}_{kh}}{\bf{w}},{\bf{v}}_{kh})_{I\times\Omega}\quad\forall{\bf{v}}_{kh}\in X^{0}_{k}(\mathbf{V}_{h});$ (5.5c)
Discrete variational inequality
$\left(\alpha\bar{{\bf{q}}}_{kh}+\bar{{\bf{z}}}_{kh},\delta{\bf{q}}-\bar{{\bf{q}}}_{kh}\right)_{I\times\Omega}\geq 0\quad\forall\delta{\bf{q}}\in\mathbf{Q}_{ad}\quad\Leftrightarrow\quad\bar{{\bf{q}}}_{kh}=P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}\left(-\frac{1}{\alpha}\bar{{\bf{z}}}_{kh}\right).$ (5.5d)
Furthermore, there exist $\bar{\mu}_{kh}^{m}\in\mathbb{R}_{\geq 0}$, $m=1,2,\dots,M$, such that the discrete Lagrange multiplier $\bar{\mu}_{kh}\in L^{1}(I)$ satisfies the expression
$\bar{\mu}_{kh}=\sum_{m=1}^{M}\frac{\bar{\mu}_{kh}^{m}}{k_{m}}\chi_{I_{m}},$ (5.6)
where $\chi_{I_{m}}$ denotes the characteristic function of the interval $I_{m}$.
Proof.
In Lemma 5.2, we have shown, that under 3 and for small enough $h$, there holds $\mathcal{G}_{kh}(\tilde{{\bf{q}}})\in\operatorname{int}(\mathcal{K}_{kh})$. Similarly to the proof of Theorem 3.7 we obtain that the optimality of $\bar{q}_{kh}$ is equivalent to the existence of a Lagrange multiplier $(\bar{\mu}_{kh}^{m})_{m=1}^{M}\in\mathbb{R}^{M}_{\geq 0}$ and the adjoint state $\bar{{\bf{z}}}_{kh}\in X^{0}_{k}(\mathbf{V}_{h})$ satisfying (5.5b), (5.5c) and (5.5d). Finally, by the construction given in (5.6), $\bar{\mu}_{kh}$ is an element of $L^{1}(I)$. ∎
Remark 5.5.
Notice, that from the definition (5.6) and using that $\bar{\mu}_{kh}\geq 0$, it holds
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\bar{\mu}_{kh}\|_{L^{1}(I)}=}\|\bar{\mu}_{kh}\|_{C(\bar{I})^{*}}=\langle\bar{\mu}_{kh},1\rangle{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\sum_{m=1}^{M}\bar{\mu}_{kh}^{m}.}$
Remark 5.6.
We would like to point out that although the state $\bar{{\bf{u}}}_{kh}$ and the adjoint $\bar{{\bf{z}}}_{kh}$ are fully discrete, the corresponding control $\bar{{\bf{q}}}_{kh}\in\mathbf{Q}_{ad}$ is piecewise constant in time via (5.5d), but not necessary piecewise polynomial in space with respect to the given mesh, due to the projection onto $[{\bf{q}}_{a},{\bf{q}}_{b}]$.
With the optimality conditions established, we now show the following stability of optimal solutions to the discrete problem subject to different discretization levels.
Lemma 5.7.
Under 3 and for $h$ small enough, there exists $C>0$ independent of $k$, $h$, such that the optimal control $\bar{{\bf{q}}}_{kh}\in\mathbf{Q}_{ad}$ of the variationally discretized problem (5.1a) - (5.1c), together with its corresponding state $\bar{{\bf{u}}}_{kh}\in X^{0}_{k}(\mathbf{V}_{h})$ and corresponding multiplier $\bar{\mu}_{kh}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{1}(\bar{I})}$ satisfy the bound
$\|\bar{\bf{q}}_{kh}\|_{L^{\infty}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{\mu}_{kh}\|_{L^{1}(\bar{I})}\leq C.$
Proof.
Due to the feasibility of $\tilde{\bf{q}}$, shown in Lemma 5.2, it holds
$J(\bar{{\bf{q}}}_{kh},\bar{{\bf{u}}}_{kh})\leq J(\tilde{{\bf{q}}},{\bf{u}}_{kh}(\tilde{{\bf{q}}}))$ $=\frac{1}{2}\lVert{\bf{u}}_{kh}(\tilde{{\bf{q}}})-{\bf{u}}_{d}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\frac{\alpha}{2}\lVert\tilde{{\bf{q}}}\rVert^{2}_{L^{2}(I)}\leq C(T,\tilde{{\bf{q}}}),$
where due to Theorem 4.2, this bound is independent of $k$ and $h$. This results in
$\|\bar{{\bf{q}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}\leq C.$ (5.7)
Let us define ${\bf{p}}=\frac{1}{2}\bar{{\bf{q}}}+\frac{1}{2}\tilde{{\bf{q}}}$. By definition ${\bf{p}}\in\mathbf{Q}_{ad}$ and thus by (5.5d) it holds $\left(\alpha\bar{{\bf{q}}}_{kh}+\bar{{\bf{z}}}_{kh},{\bf{p}}-\bar{{\bf{q}}}_{kh}\right)_{I\times\Omega}\geq 0.$ This yields
00 0 $\leq\alpha\left(\bar{{\bf{q}}}_{kh},{\bf{p}}-\bar{{\bf{q}}}_{kh}\right)_{I\times\Omega}+\left(\bar{{\bf{z}}}_{kh},{\bf{p}}-\bar{{\bf{q}}}_{kh}\right)_{I\times\Omega}$ (5.8)
$=\alpha\left(\bar{{\bf{q}}}_{kh},{\bf{p}}-\bar{{\bf{q}}}_{kh}\right)_{I\times\Omega}+B({\bf{u}}_{kh}({\bf{p}})-\bar{{\bf{u}}}_{kh},\bar{{\bf{z}}}_{kh})$
$=\alpha\left(\bar{{\bf{q}}}_{kh},{\bf{p}}-\bar{{\bf{q}}}_{kh}\right)_{I\times\Omega}+({\bf{u}}_{kh}({\bf{p}})-\bar{{\bf{u}}}_{kh},\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d})_{I\times\Omega}+\langle\bar{\mu}_{kh},G_{\bf{w}}({{\bf{u}}}_{kh}({\bf{p}}))-G_{\bf{w}}(\bar{{\bf{u}}}_{kh})\rangle,$
where we can bound the first two terms by (5.7) and obtain
$0\leq C+\langle\bar{\mu}_{kh},G_{\bf{w}}({{\bf{u}}}_{kh}({\bf{p}}))-G_{\bf{w}}(\bar{{\bf{u}}}_{kh})\rangle.$ (5.9)
For ${\bf{p}}$, using 3, we can follow a similar argument as in the proof of Lemma 5.2, in order to obtain
$G_{\bf w}({\bf{u}}_{kh}({{\bf{p}}}))$ $=G_{\bf w}({\bf{u}}_{kh}({{\bf{p}}})-{\bf{u}}({{\bf{p}}}))+G_{\bf w}({\bf{u}}({{\bf{p}}}))$
$=G_{\bf w}({\bf{u}}_{kh}({{\bf{p}}})-{\bf{u}}({{\bf{p}}}))+\frac{1}{2}G_{\bf w}(\bar{{\bf{u}}})+\frac{1}{2}G_{\bf w}({\bf{u}}(\tilde{{\bf{q}}}))$
$\leq\|{\bf{w}}\|_{L^{2}(\Omega)}\|{\bf{u}}_{kh}({{\bf{p}}})-{\bf{u}}({{\bf{p}}})\|_{L^{\infty}(I;L^{2}(\Omega))}+\frac{1}{2}G_{\bf w}(\bar{{\bf{u}}})+\frac{1}{2}G_{\bf w}({\bf{u}}(\tilde{{\bf{q}}}))$
$\leq\frac{1}{4}\delta+\frac{1}{2}\beta+\frac{1}{2}\beta-\frac{1}{2}\delta=\beta-\frac{1}{4}\delta.$
Inserting this into (5.9), yields together with $\bar{\mu}_{kh}\geq 0$ and the complementarity conditions (5.5b):
$0\leq C+\langle\bar{\mu}_{kh},b-G_{\bf{w}}(\bar{{\bf{u}}}_{kh})\rangle-\frac{\delta}{4}\langle\bar{\mu}_{kh},1\rangle=C-\frac{\delta}{4}\langle\bar{\mu}_{kh},1\rangle.$
Thus, again using $\bar{\mu}_{kh}\geq 0$ and Remark 5.5, results in
$\|\bar{\mu}_{kh}\|_{L^{1}(\bar{I})}=\langle\bar{\mu}_{kh},1\rangle\leq C.$ (5.10)
Combining (5.7) and (5.10) with Remark 4.3 yields the boundedness of $\bar{\bf{z}}_{kh}$ in $L^{\infty}(I;L^{2}(\Omega)^{d})$. By the representation $\bar{\bf{q}}_{kh}=P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}\left(-\frac{1}{\alpha}\bar{\bf{z}}_{kh}\right)$, this shows $\|\bar{\bf{q}}_{kh}\|_{L^{\infty}(I;L^{2}(\Omega))}\leq C$, which concludes the proof. ∎
Theorem 5.8.
Let 3 hold and let $h$ be sufficiently small. Let $(\bar{\bf{q}},\bar{\bf{u}})$ and $(\bar{\bf{q}}_{kh},\bar{\bf{u}}_{kh})$ be the unique solutions to the continuous and variationally discretized optimal control problems (1.1a) -(1.1d) and (5.1a)-(5.1c) respectively. Then there exists a constant $C>0$, such that it holds
$\sqrt{\alpha}\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{kh}\rVert_{L^{2}(I;L^{2}(\Omega))}+\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{kh}\rVert_{L^{2}(I;L^{2}(\Omega))}\leq C\,\ell_{k}(k^{\frac{1}{2}}+h),\quad\ell_{k}=\ln{\frac{T}{k}}.$
Proof.
Choosing $\delta{\bf{q}}=\bar{{\bf{q}}}_{kh}$ in (3.6d) and $\delta{\bf{q}}=\bar{{\bf{q}}}$ in (5.5d), results in
$\left(\alpha\bar{{\bf{q}}}+\bar{{\bf{z}}},\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}\right)_{I\times\Omega}\geq 0\quad\text{and}\quad\left(\alpha\bar{{\bf{q}}}_{kh}+\bar{{\bf{z}}}_{kh},\bar{{\bf{q}}}-\bar{{\bf{q}}}_{kh}\right)_{I\times\Omega}\geq 0.$ (5.11)
Adding these two inequalities results in
$\alpha\|\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}\|^{2}_{L^{2}(I;L^{2}(\Omega))}\leq(\bar{{\bf{z}}},\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}})_{I\times\Omega}+(\bar{{\bf{z}}}_{kh},\bar{{\bf{q}}}-\bar{{\bf{q}}}_{kh})_{I\times\Omega}:=I_{1}+I_{2}.$ (5.12)
We estimate the two terms separately.
Estimate for $I_{1}$. Using the weak formulations (2.4) and (3.2) of the continuous state and adjoint equations (3.6a) & (3.6c), respectively, we have
$I_{1}$ $=\left\langle\partial_{t}({\bf{u}}(\bar{{\bf{q}}}_{kh})-\bar{{\bf{u}}}),\bar{\bf{z}}\right\rangle_{I\times\Omega}+\left({\rm NewA}({\bf{u}}(\bar{{\bf{q}}}_{kh})-\bar{{\bf{u}}}),{\rm NewA}\bar{\bf{z}}\right)_{I\times\Omega}$
$=(\bar{{\bf{u}}}-{\bf{u}}_{d},{\bf{u}}(\bar{{\bf{q}}}_{kh})-\bar{{\bf{u}}})_{I\times\Omega}+\langle G_{\bf w}({\bf{u}}(\bar{{\bf{q}}}_{kh})-\bar{{\bf{u}}}),\bar{\mu}\rangle.$
Introducing the pointwise projection onto the state constraint
$P_{\beta}v=P_{\beta}v(t):=\max\{v(t),\beta\},\quad\forall t\in\bar{I},$ (5.13)
the last term can be estimated as
$\langle G_{\bf w}({\bf{u}}(\bar{{\bf{q}}}_{kh})-\bar{{\bf{u}}}),\bar{\mu}\rangle$ $=\langle G_{\bf w}({\bf{u}}(\bar{{\bf{q}}}_{kh}))-P_{\beta}G_{\bf w}({\bf{u}}(\bar{{\bf{q}}}_{kh})),\bar{\mu}\rangle+\langle P_{\beta}G_{\bf w}({\bf{u}}(\bar{{\bf{q}}}_{kh}))-G_{\bf w}(\bar{{\bf{u}}}),\bar{\mu}\rangle$
$\leq\langle G_{\bf w}({\bf{u}}(\bar{{\bf{q}}}_{kh}))-P_{\beta}G_{\bf w}({\bf{u}}(\bar{{\bf{q}}}_{kh})),\bar{\mu}\rangle+\langle{\beta}-G_{\bf w}(\bar{{\bf{u}}}),\bar{\mu}\rangle,$
where we have used, that due to $\bar{\mu}\geq 0$, it holds $\langle P_{\beta}G_{\bf{w}}({\bf{u}}(\bar{\bf{q}}_{kh})),\bar{\mu}\rangle\leq\langle{\beta},\bar{\mu}\rangle$. Using the complementarity condition (3.6b), it holds $\langle{\beta}-G_{\bf w}(\bar{{\bf{u}}}),\bar{\mu}\rangle=0$, hence we have
$\langle G_{\bf w}({\bf{u}}(\bar{{\bf{q}}}_{kh})-\bar{{\bf{u}}}),\bar{\mu}\rangle\leq\langle G_{\bf w}({\bf{u}}(\bar{{\bf{q}}}_{kh}))-P_{\beta}G_{\bf w}({\bf{u}}(\bar{{\bf{q}}}_{kh})),\bar{\mu}\rangle.$
Now using that
$|P_{\beta}v-P_{\beta}u|\leq|v-u|,$ (5.14)
by the triangle inequality and using that $G_{\bf w}(\bar{{\bf{u}}}_{kh})\leq{\beta}$, we obtain
$\langle G_{\bf w}({\bf{u}}(\bar{{\bf{q}}}_{kh}))-P_{\beta}G_{\bf w}({\bf{u}}(\bar{{\bf{q}}}_{kh})),\bar{\mu}\rangle\leq$ $\|G_{\bf w}({\bf{u}}(\bar{{\bf{q}}}_{kh}))-P_{\beta}G_{\bf w}({\bf{u}}(\bar{{\bf{q}}}_{kh})\|_{L^{\infty}(I)}\|\bar{\mu}\|_{C(\bar{I})^{*}}$
$\leq$ $\|G_{\bf w}({\bf{u}}(\bar{{\bf{q}}}_{kh}))-G_{\bf w}(\bar{{\bf{u}}}_{kh})\|_{L^{\infty}(I)}\|\bar{\mu}\|_{C(\bar{I})^{*}}$
$+\|P_{\beta}G_{\bf w}(\bar{{\bf{u}}}_{kh})-P_{\beta}G_{\bf w}({\bf{u}}(\bar{{\bf{q}}}_{kh})\|_{L^{\infty}(I)}\|\bar{\mu}\|_{C(\bar{I})^{*}}$
$\leq$ $2\|{\bf w}\|_{L^{2}(\Omega)}\|\bar{{\bf{u}}}_{kh}(\bar{{\bf{q}}}_{kh})-{\bf{u}}(\bar{{\bf{q}}}_{kh})\|_{L^{\infty}(I;L^{2}(\Omega))}\|\bar{\mu}\|_{C(\bar{I})^{*}}$
$\leq$ $C\ell_{k}^{2}(h^{2}+k)\|\bar{{\bf{q}}}_{kh}\|_{L^{\infty}(I;L^{2}(\Omega))}\|{\bf w}\|_{L^{2}(\Omega)}\|\bar{\mu}\|_{C(\bar{I})^{*}}$
$\leq$ $C\ell_{k}^{2}(h^{2}+k),$
where in the last two steps we used Corollary 4.5 and Lemma 5.7. Thus,
$I_{1}\leq C\ell_{k}^{2}(h^{2}+k)+(\bar{{\bf{u}}}-{\bf{u}}_{d},{\bf{u}}(\bar{{\bf{q}}}_{kh})-\bar{{\bf{u}}})_{I\times\Omega}.$
Estimate for $I_{2}$. Similarly, using the fully discrete state and adjoint equations (5.5a) and (5.5c), respectively, we have
$I_{2}$ $=B({\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}}_{kh},\bar{{\bf{z}}}_{kh})$
$=(\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d},{\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}}_{kh})_{I\times\Omega}+\langle G_{\bf w}({\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}}_{kh}),\bar{\mu}_{kh}\rangle.$
Using the projection $P_{\beta}$ defined in (5.13), the last term in $I_{2}$ can be estimated as
$\langle G_{\bf w}({\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}}_{kh}),\bar{\mu}\rangle$ $=\langle G_{\bf w}({\bf{u}}_{kh}(\bar{{\bf{q}}}))-P_{\beta}G_{\bf w}({\bf{u}}_{kh}(\bar{{\bf{q}}})),\bar{\mu}_{kh}\rangle+\langle P_{\beta}G_{\bf w}({\bf{u}}_{kh}(\bar{{\bf{q}}}))-G_{\bf w}(\bar{{\bf{u}}}_{kh}),\bar{\mu}_{kh}\rangle$
$\leq\langle G_{\bf w}({\bf{u}}_{kh}(\bar{{\bf{q}}}))-P_{\beta}G_{\bf w}({\bf{u}}_{kh}(\bar{{\bf{q}}})),\bar{\mu}_{kh}\rangle+\langle{\beta}-G_{\bf w}(\bar{{\bf{u}}}_{kh}),\bar{\mu}_{kh}\rangle,$
where we have used, that due to $\bar{\mu}_{kh}\geq 0$, it holds $\langle P_{\beta}G_{\bf{w}}({\bf{u}}_{kh}(\bar{\bf{q}})),\bar{\mu}_{kh}\rangle\leq\langle{\beta},\bar{\mu}_{kh}\rangle$. Using the complementarity condition (5.5b), it holds $\langle{\beta}-G_{\bf{w}}(\bar{{\bf{u}}}_{kh}),\bar{\mu}_{kh}\rangle=0$, hence we have
$\langle G_{\bf w}({\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}}_{kh}),\bar{\mu}\rangle\leq\langle G_{\bf w}({\bf{u}}_{kh}(\bar{{\bf{q}}}))-P_{\beta}G_{\bf w}({\bf{u}}_{kh}(\bar{{\bf{q}}})),\bar{\mu}_{kh}\rangle.$
Using (5.14), the triangle inequality and using that $G_{\bf w}(\bar{{\bf{u}}}_{kh})\leq{\beta}$, we obtain
$\langle G_{\bf w}({\bf{u}}_{kh}(\bar{{\bf{q}}}))-P_{\beta}G_{\bf w}({\bf{u}}_{kh}(\bar{{\bf{q}}}),\bar{\mu}_{kh}\rangle\leq$ $\|G_{\bf w}({\bf{u}}_{kh}(\bar{{\bf{q}}}))-P_{\beta}G_{\bf w}({\bf{u}}_{kh}(\bar{{\bf{q}}})\|_{L^{\infty}(I)}\|\bar{\mu}_{kh}\|_{L^{1}(\bar{I})}$
$\leq$ $\|G_{\bf w}({\bf{u}}_{kh}(\bar{{\bf{q}}}))-G_{\bf w}(\bar{{\bf{u}}})\|_{L^{\infty}(I)}\|\bar{\mu}_{kh}\|_{L^{1}(\bar{I})}$
$+\|P_{\beta}G_{\bf w}(\bar{{\bf{u}}})-P_{\beta}G_{\bf w}({\bf{u}}_{kh}(\bar{{\bf{q}}})\|_{L^{\infty}(I)}\|\bar{\mu}_{kh}\|_{L^{1}(\bar{I})}$
$\leq$ $2\|{\bf w}\|_{L^{2}(\Omega)}\|\bar{{\bf{u}}}(\bar{{\bf{q}}})-({\bf{u}}_{kh}(\bar{{\bf{q}}})\|_{L^{\infty}(I;L^{2}(\Omega))}\|\bar{\mu}_{kh}\|_{L^{1}(\bar{I})}$
$\leq$ $C\ell_{k}^{2}(h^{2}+k)\|\bar{{\bf{q}}}\|_{L^{\infty}(I;L^{2}(\Omega))}\|{\bf w}\|_{L^{2}(\Omega)}\|\bar{\mu}_{kh}\|_{L^{1}(\bar{I})}$
$\leq$ $C\ell_{k}^{2}(h^{2}+k),$
where in the last two steps we used Corollary 4.5 and Lemma 5.7. Thus,
$I_{2}\leq C\ell_{k}^{2}(h^{2}+k)+(\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d},{\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}}_{kh})_{I\times\Omega}.$
Combining the estimates for $I_{1}$ and $I_{2}$ and using that
$(\bar{{\bf{u}}}-{\bf{u}}_{d},{\bf{u}}(\bar{{\bf{q}}}_{kh})$ $-\bar{{\bf{u}}})_{I\times\Omega}+(\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d},{\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}}_{kh})_{I\times\Omega}$
$=$ $(\bar{{\bf{u}}}-{\bf{u}}_{d},\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}})_{I\times\Omega}+(\bar{{\bf{u}}}-{\bf{u}}_{d},{\bf{u}}(\bar{{\bf{q}}}_{kh})-\bar{{\bf{u}}}_{kh})_{I\times\Omega}$
$+(\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d},{\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}})_{I\times\Omega}+(\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d},\bar{{\bf{u}}}-\bar{{\bf{u}}}_{kh})_{I\times\Omega}$
$=$ $-\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{kh}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+(\bar{{\bf{u}}}-{\bf{u}}_{d},{\bf{u}}(\bar{{\bf{q}}}_{kh})-\bar{{\bf{u}}}_{kh})_{I\times\Omega}+(\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d},{\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}})_{I\times\Omega}$
by using Corollary 4.7, we obtain
$\alpha$ $\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{kh}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{kh}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}$
$\leq C\ell^{2}_{k}(k+h^{2})+(\bar{{\bf{u}}}-{\bf{u}}_{d},{\bf{u}}(\bar{{\bf{q}}}_{kh})-\bar{{\bf{u}}}_{kh})_{I\times\Omega}+(\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d},{\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}})_{I\times\Omega}$
$\leq C\ell^{2}_{k}(k+h^{2})+\|\bar{{\bf{u}}}-{\bf{u}}_{d}\|_{L^{2}(I;L^{2}(\Omega))}\|{\bf{u}}(\bar{{\bf{q}}}_{kh})-\bar{{\bf{u}}}_{kh}(\bar{{\bf{q}}}_{kh})\|_{L^{2}(I;L^{2}(\Omega))}$
$\quad+\|\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d}\|_{L^{2}(I;L^{2}(\Omega))}\|{\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}}(\bar{{\bf{q}}})\|_{L^{2}(I;L^{2}(\Omega))}$
$\leq C\ell^{2}_{k}(k+h^{2})\left[1+\left(\|\bar{{\bf{u}}}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}+\|{\bf{u}}_{d}\|_{L^{2}(I;L^{2}(\Omega))}\right)\left(\|\bar{{\bf{q}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{{\bf{q}}}\|_{L^{2}(I;L^{2}(\Omega))}\right)\right]$
$\leq C\ell^{2}_{k}(k+h^{2}),$
where in the last step, we used the boundedness of $\|\bar{{\bf{u}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}$, $\|\bar{{\bf{u}}}\|_{L^{2}(I;L^{2}(\Omega))}$, $\|\bar{{\bf{q}}}\|_{L^{2}(I;L^{2}(\Omega))}$, and $\|\bar{{\bf{q}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}$ from Theorem 3.8 and Lemma 5.7. ∎
6 Full Discretization of the Optimal Control Problem
We discretize the control by piecewise constant functions on the same partition as the fully discrete approximation of the state and adjoint variables. We set
$\mathbf{Q}_{0}=\{{\bf{q}}\in L^{2}(I;L^{2}(\Omega)^{d}):\ {\bf{q}}\mid_{I_{m}\times K}\in\mathbb{P}_{0}(I_{m};\mathbb{P}_{0}(K)^{d}),\ m=1,2,\dots,M,\ K\in\mathcal{T}_{h}\}.$ (6.1)
We also define the corresponding admissible set
$\mathbf{Q}_{0,ad}:=\mathbf{Q}_{0}\cap\mathbf{Q}_{ad}.$
We introduce the projection $\pi_{d}\colon L^{2}(I;L^{2}(\Omega)^{d})\to\mathbf{Q}_{0}$, defined by
$\left({\bf{q}}-\pi_{d}{\bf{q}},\mathbf{r}\right)_{I\times\Omega}=0\qquad\forall\mathbf{r}\in\mathbf{Q}_{0}.$ (6.2)
Note that this projection satisfies the explicit formula
$\pi_{d}{\bf{q}}|_{I_{m}\times K}=\frac{1}{k_{m}|K|}\int_{I_{m}\times K}{\bf{q}}(t,x)\ d(t,x)\qquad\text{for all }m=1,...,M,\ K\in\mathcal{T}_{h}.$
Hence it is straightforward to check, that this $L^{2}$ projection onto piecewise constants is stable in $L^{\infty}(I\times\Omega)^{d}$ and $L^{\infty}(I;L^{2}(\Omega)^{d})$ and there holds
$\|\pi_{d}{\bf{q}}\|_{L^{\infty}(I\times\Omega)}\leq\|{\bf{q}}\|_{L^{\infty}(I\times\Omega)}\quad\text{and}\quad\|\pi_{d}{\bf{q}}\|_{L^{\infty}(I;L^{2}(\Omega))}\leq\|{\bf{q}}\|_{L^{\infty}(I;L^{2}(\Omega))}.$ (6.3)
Further, we have $\pi_{d}(\mathbf{Q}_{ad})\subset\mathbf{Q}_{0,ad}$. We can now formulate the fully discrete optimal control problem, which reads
Minimize $\text{Minimize }\;J({\bf{q}}_{\sigma},{\bf{u}}_{\sigma})=\frac{1}{2}\lVert{\bf{u}}_{\sigma}-{\bf{u}}_{d}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\frac{\alpha}{2}\lVert{\bf{q}}_{\sigma}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}$ (6.4a)
subject to $({\bf{q}}_{\sigma},{\bf{u}}_{\sigma})\in\mathbf{Q}_{0,ad}\times X^{0}_{k}(\mathbf{V}_{h})$, satisfying
$B({\bf{u}}_{\sigma},{\bf{v}}_{kh})=({\bf{q}}_{\sigma},{\bf{v}}_{kh})_{I\times\Omega}\quad\text{for all}\quad{\bf{v}}_{kh}\in X^{0}_{k}(\mathbf{V}_{h})$ (6.4b)
and
$G_{\bf{w}}({\bf{u}}_{\sigma})\mid_{I_{m}}\leq\beta\quad\text{for}\quad m=1,2,\dots,M.$ (6.4c)
The following lemma guarantees that also for the fully discrete optimal control problem, there exist feasible controls such that the associated fully discrete state strictly satisfies the state constraint.
Lemma 6.1.
Let 3 be satisfied and let $h$ be sufficiently small. Then the projection $\pi_{d}\tilde{\bf{q}}\in\mathbf{Q}_{0,ad}$ of the Slater point $\tilde{\bf{q}}\in\mathbf{Q}_{ad}$ from 1 satisfies the following discrete Slater condition
$G_{\bf{w}}({\bf{u}}_{kh}(\pi_{d}\tilde{q}))<\beta{\qquad\text{for all }t\in\bar{I}}.$
Proof.
From Lemma 5.2, we know, that there exists $\delta>0$ such that $G_{\bf{w}}({\bf{u}}_{kh}(\tilde{q}))<\beta-\delta$. As the discrete solution operator $S_{kh}$ is linear and continuous from $L^{2}(I;L^{2}(\Omega)^{d})\to L^{\infty}(I;L^{2}(\Omega)^{d})$, we have
$G_{\bf{w}}({\bf{u}}_{kh}(\pi_{d}\tilde{\bf{q}}))=G_{\bf{w}}({\bf{u}}_{kh}(\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}))+G_{\bf{w}}({\bf{u}}_{kh}(\tilde{\bf{q}}))<C\|\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}\|_{L^{2}(I;L^{2}(\Omega))}+\beta-\delta.$
As 3 guarantees $k\to 0$ as $h\to 0$, it holds $\|\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}\|_{L^{2}(I;L^{2}(\Omega))}\to 0$ for $h\to 0$. This implies that for $h$ small enough, we have $C\|\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}\|_{L^{2}(I;L^{2}(\Omega))}\leq\delta$ and as a consequence $G_{\bf{w}}({\bf{u}}_{kh}(\pi_{d}\tilde{\bf{q}}))<\beta$. ∎
Theorem 6.2.
Let 3 be satisfied and let $h$ be sufficiently small. Then there exists a unique solution $(\bar{\bf{q}}_{\sigma},\bar{\bf{u}}_{\sigma})$ to the fully discrete optimal control problem (6.4a)-(6.4c).
Proof.
As Lemma 6.1 shows feasibility of $(\pi_{d}\tilde{\bf{q}},{\bf{u}}_{kh}(\pi_{d}\tilde{\bf{q}}))$ under the given assumptions, the existence proof follows the same steps as the one of Theorem 3.6 on the continuous level. ∎
Similar to Section 5, we can rewrite the problem (6.4a)-(6.4c) in the reduced form
$\text{Minimize }\ j_{kh}({\bf{q}}_{\sigma})=J({\bf{q}}_{\sigma},S_{kh}({\bf{q}}_{\sigma}))\quad\text{over}\quad{\bf{q}}_{\sigma}\in\mathbf{Q}_{0,ad}\quad\text{subject to}\quad\mathcal{G}_{kh}({\bf{q}}_{\sigma})\in\mathcal{K}_{kh}.$ (6.5)
Note that compared to the the variationally discretized optimal control problem (5.3), only the control space has changed.
Theorem 6.3 (First order optimality conditions for discretized controls).
A control $\bar{{\bf{q}}}_{\sigma}\in\mathbf{Q}_{0,ad}$ and the associated state $\bar{{\bf{u}}}_{\sigma}={\bf{u}}_{kh}(\bar{{\bf{q}}}_{\sigma})\in X^{0}_{k}(\mathbf{V}_{h})$ is the optimal solution to the problem (6.4a)-(6.4c) if and only if there exists an adjoint state $\bar{{\bf{z}}}_{\sigma}\in X^{0}_{k}(\mathbf{V}_{h})$ and a Lagrange multiplier $\bar{\mu}_{\sigma}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{1}(I)}$ that satisfy:
Discrete state equation
$B(\bar{{\bf{u}}}_{\sigma},{\bf{v}}_{kh})=(\bar{{\bf{q}}}_{\sigma},{\bf{v}}_{kh})_{I\times\Omega},\quad\forall{\bf{v}}_{kh}\in X^{0}_{k}(\mathbf{V}_{h});$ (6.6a)
Discrete state constraint and complementarity conditions
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}G_{\bf{w}}(\bar{\bf{u}}_{\sigma})|_{I_{m}}\leq\beta,\ m=1,...,M,\quad}\bar{\mu}_{\sigma}\geq 0\quad\text{and}\quad\langle\bar{\mu}_{\sigma},\beta-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}G_{\bf{w}}(\bar{\bf{u}}_{\sigma})}\rangle=0;$ (6.6b)
Discrete adjoint equation
$B({\bf{v}}_{kh},\bar{{\bf{z}}}_{\sigma})=(\bar{{\bf{u}}}_{\sigma}-{\bf{u}}_{d}+{\bar{\mu}_{\sigma}}{\bf{w}},{\bf{v}}_{kh})_{I\times\Omega}\quad\forall{\bf{v}}_{kh}\in X^{0}_{k}(\mathbf{V}_{h});$ (6.6c)
Discrete variational inequality
$\left(\alpha\bar{{\bf{q}}}_{\sigma}+\bar{{\bf{z}}}_{\sigma},\delta{\bf{q}}-\bar{{\bf{q}}}_{\sigma}\right)_{I\times\Omega}\geq 0\quad\forall\delta{\bf{q}}\in\mathbf{Q}_{0,ad}\quad\Leftrightarrow\quad\bar{{\bf{q}}}_{\sigma}=P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}\left(-\frac{1}{\alpha}\bar{{\bf{z}}}_{\sigma}\right).$ (6.6d)
Furthermore, there exist $\bar{\mu}_{\sigma}^{m}\in\mathbb{R}_{\geq 0}$, $m=1,2,\dots,M$, such that the discrete Lagrange multiplier $\bar{\mu}_{\sigma}\in L^{1}(I)$ satisfies the expression
$\bar{\mu}_{\sigma}=\sum_{m=1}^{M}\frac{\bar{\mu}_{\sigma}^{m}}{k_{m}}\chi_{I_{m}},$ (6.7)
where $\chi_{I_{m}}$ denotes the characteristic function of the interval $I_{m}$.
Proof.
The proof is almost identical to the proof of Theorem 5.4. ∎
Again, due to $\bar{\mu}_{\sigma}\geq 0$, it holds $\|\bar{\mu}_{\sigma}\|_{L^{1}(I)}=\|\bar{\mu}_{\sigma}\|_{(C(\bar{I})^{*})}=\langle\bar{\mu}_{\sigma},1\rangle=\sum_{m=1}^{M}\bar{\mu}_{\sigma}^{m}$.
Lemma 6.4.
Let 3 be satisfied, and $h$ be small enough. Then there exists a constant $C>0$ independent of $k$ and $h$, such that the fully discrete optimal control $\bar{{\bf{q}}}_{\sigma}\in\mathbf{Q}_{0,ad}$, solving (6.4a)-(6.4c), together with its corresponding state $\bar{{\bf{u}}}_{\sigma}\in X^{0}_{k}(\mathbf{V}_{h})$ and corresponding multiplier $\bar{\mu}_{\sigma}\in L^{1}(\bar{I})$ satisfies the bound
$\|\bar{{\bf{q}}}_{\sigma}\|_{L^{\infty}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{\mu}_{\sigma}\|_{(C(\bar{I}))^{*}}\leq C.$
Proof.
By Lemma 6.1 under the given assumptions, the fully discrete control $\pi_{d}\tilde{\bf{q}}$ is feasible, and thus it holds
$J(\bar{{\bf{q}}}_{\sigma},\bar{{\bf{u}}}_{\sigma})$ $\leq J(\pi_{d}\tilde{{\bf{q}}},\bar{{\bf{u}}}_{kh}(\pi_{d}\tilde{{\bf{q}}}))$
$=\frac{1}{2}\lVert\bar{{\bf{u}}}_{kh}(\pi_{d}\tilde{{\bf{q}}})-{\bf{u}}_{d}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\frac{\alpha}{2}\lVert\pi_{d}\tilde{{\bf{q}}}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}$
$\leq\lVert\bar{{\bf{u}}}_{kh}(\tilde{{\bf{q}}})-{\bf{u}}(\tilde{{\bf{q}}})\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\lVert{\bf{u}}(\tilde{{\bf{q}}})-{\bf{u}}_{d}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\frac{\alpha}{2}\lVert\tilde{{\bf{q}}}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}$
$\leq C(T,\tilde{{\bf{q}}},{\bf{u}}_{d}).$
As a result we obtain $\|\bar{{\bf{q}}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}\leq C.$ The proof of $\|\bar{\mu}_{\sigma}\|_{C(\bar{I})^{*})}\leq C$ and $\|\bar{\bf{q}}_{\sigma}\|_{L^{\infty}(I;L^{2}(\Omega))}\leq C$ is then accomplished by following the same steps as the proof of Lemma 5.7, making use of $G_{\bf{w}}({\bf{u}}_{kh}({\pi_{d}{\bf{p}}}))\leq\beta-\frac{1}{4}\delta$ for $h$ sufficiently small.
Theorem 6.5.
Let 3 be satisfied and let $h$ be small enough. Let $(\bar{\bf{q}}_{kh},\bar{\bf{u}}_{kh})$ and $(\bar{\bf{q}}_{\sigma},\bar{\bf{u}}_{\sigma})$ denote the optimal solutions of the variationally discretized optimal control problem (5.1a)-(5.1c) and the fully discretized optimal control problem (6.4a)-(6.4c). Then there exists a constant $C>0$ such that it holds
$\alpha\lVert\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}_{\sigma}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\lVert\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{\sigma}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}\leq Ch^{2}.$
Proof.
Choosing $\delta{\bf{q}}=\bar{{\bf{q}}}_{\sigma}\in\mathbf{Q}_{ad}$ in (5.5d) and $\delta{\bf{q}}=\pi_{d}\bar{{\bf{q}}}_{kh}\in\mathbf{Q}_{0,ad}$ in (6.6d), results in
$\left(\alpha\bar{{\bf{q}}}_{kh}+\bar{{\bf{z}}}_{kh},\bar{{\bf{q}}}_{\sigma}-\bar{{\bf{q}}}_{kh}\right)_{I\times\Omega}\geq 0\quad\text{and}\quad\left(-\alpha\bar{{\bf{q}}}_{\sigma}-\bar{{\bf{z}}}_{\sigma},\bar{{\bf{q}}}_{\sigma}-\pi_{d}\bar{{\bf{q}}}_{kh}\right)_{I\times\Omega}\geq 0.$ (6.8)
Adding these two inequalities, we obtain
$\alpha\|\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}_{\sigma}\|^{2}_{L^{2}(I;L^{2}(\Omega))}\leq(\bar{{\bf{z}}}_{kh}-\bar{{\bf{z}}}_{\sigma},\bar{{\bf{q}}}_{\sigma}-\bar{{\bf{q}}}_{kh})_{I\times\Omega}+(\alpha\bar{{\bf{q}}}_{\sigma}+\bar{{\bf{z}}}_{\sigma},\pi_{d}\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}_{kh})_{I\times\Omega}:=I_{1}+I_{2}.$ (6.9)
We estimate the two terms separately.
Estimate for $I_{1}$. Using the discrete state equations (5.5a),(6.6a) the corresponding adjoint equations (5.5c) and (6.6c), respectively, we have
$I_{1}$ $=B(\bar{{\bf{u}}}_{\sigma}-\bar{{\bf{u}}}_{kh},\bar{{\bf{z}}}_{kh}-\bar{{\bf{z}}}_{\sigma})$
$=(\bar{{\bf{u}}}_{\sigma}-\bar{{\bf{u}}}_{kh},\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{\sigma})_{I\times\Omega}+\langle G_{\bf w}(\bar{{\bf{u}}}_{\sigma})-G_{\bf w}(\bar{{\bf{u}}}_{kh}),\bar{\mu}_{kh}\rangle-\langle G_{\bf w}(\bar{{\bf{u}}}_{\sigma})-G_{\bf w}(\bar{{\bf{u}}}_{kh}),\bar{\mu}_{\sigma}\rangle$
$\leq-\lVert\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{\sigma}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\langle\beta-G_{\bf w}(\bar{{\bf{u}}}_{kh}),\bar{\mu}_{kh}\rangle+\langle\beta-G_{\bf w}(\bar{{\bf{u}}}_{\sigma}),\bar{\mu}_{\sigma}\rangle$
$=-\lVert\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{\sigma}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}.$
By the Cauchy-Schwarz inequality and properties of the $L^{2}$-projection
$I_{2}$ $=\left(\alpha\bar{\bf{q}}_{\sigma}+\bar{\bf{z}}_{\sigma},\pi_{d}\bar{\bf{q}}_{kh}-\bar{\bf{q}}_{kh}\right)_{I\times\Omega}$
$=\left(\bar{\bf{z}}_{\sigma}-\pi_{d}\bar{\bf{z}}_{\sigma},\pi_{d}\bar{\bf{q}}_{kh}-\bar{\bf{q}}_{kh}\right)_{I\times\Omega}$
$\leq Ch^{2}\|\nabla\bar{\bf{z}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}\|\nabla\bar{\bf{q}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}.$
Using that $\|{\rm NewA}\bar{{\bf{q}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}\leq\alpha^{-1}\|{\rm NewA}\bar{{\bf{z}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}$, the stability of $P_{ad}$ and the stability of solutions to the fully discrete dual problem, pointed out in Remark 4.3, gives
$I_{2}\leq Ch^{2}\left(\|\bar{{\bf{u}}}_{\sigma}-\bar{{\bf{u}}}_{d}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{\mu}_{\sigma}\|_{L^{1}(I)}\|{\bf{w}}\|_{L^{2}(\Omega)}\right)\left(\|\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{d}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{\mu}_{kh}\|_{L^{1}(I)}\|{\bf{w}}\|_{L^{2}(\Omega)}\right).$
Now the boundedness of $\|\bar{{\bf{u}}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}$, $\|\bar{{\bf{u}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}$, $\|\bar{\mu}_{\sigma}\|_{L^{1}(I)}$, and $\|\bar{\mu}_{kh}\|_{L^{1}(I)}$ from Lemmas 5.7 and 6.4 finish the proof. ∎
With this last error estimate, our main result now directly follows from Theorem 5.8 and Theorem 6.5.
Theorem 6.6 (Error estimate for the control).
Let 3 be satisfied and let $h$ be small enough. Let $\bar{\bf{q}}\in\mathbf{Q}_{ad}$ and $\bar{\bf{q}}_{\sigma}\in\mathbf{Q}_{0,ad}$ be the solutions to the continuous and fully discrete optimal control problems (1.1a)-(1.1d) and (6.4a)-(6.4c).
$\sqrt{\alpha}\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{\sigma}\rVert_{L^{2}(I;L^{2}(\Omega))}+\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{\sigma}\rVert_{L^{2}(I;L^{2}(\Omega))}\leq C\,\ell_{k}(k^{\frac{1}{2}}+h),\quad\ell_{k}=\ln{\frac{T}{k}}.$
Remark 6.7.
All of our main results presented in Theorems 5.8, 6.5 and 6.6 do not require any additional regularity assumptions on the optimal control $\bar{\bf{q}}$, but work precisely with the regularity that is obtained from the optimality conditions. Furthermore, the techniques used in proving these results allow us to avoid strong coupling conditions on $k$ and $h$. We only require the product $\ell_{k}h$ to converge to zero, in order to obtain a Slater condition for the discrete problems, which can be guaranteed by a very mild coupling condition as in 3.
7 Improved regularity
Despite having no a priori smoothness regularity for $\bar{{\bf{z}}}$ and as a result for $\bar{{\bf{q}}}$, our main result shows almost $k^{\frac{1}{2}}$ convergence rate for the error of the optimal control. Similarly to [29], we can use this result to establish improved regularity for the optimal control.
Theorem 7.1.
Let $\bar{{\bf{q}}}\in\mathbf{Q}_{ad}$ be the optimal solution to (1.1a)-(1.1d). Then,
$\bar{{\bf{q}}}\in L^{2}(I;H^{1}(\Omega)^{d})\cap H^{s}(I;L^{2}(\Omega)^{d}),\quad\forall s<\frac{1}{2}.$
Additionally, if ${\bf{q}}_{a},{\bf{q}}_{b}\in\mathbb{R}^{d}$, i.e., are finite, it holds $\bar{\bf{q}}\in{L^{\infty}(I\times\Omega)^{d}}$.
Proof.
The proof is identical to the proof of Theorem 7.1 in [29]. ∎
In light of Theorem 3.10 and the imbedding $\mathrm{BV}(I)\hookrightarrow H^{s}(I)$ for all $s<\frac{1}{2}$, this result is particularly interesting, as it holds without additional regularity assumptions on the weight ${\bf{w}}$. Note also, that the regulartiy properties of Theorem 3.10 and its corrollaries were not used in the derivation of our error estimates. It remains an open question, how these results can be used to derive improved error estimates for the optimal control problem directly.
8 Numerical results
In the following, we present two numerical examples, studying the orders of convergence for the optimal control problem. We first present an example for smooth data. These results can be compared to the numerical example in [29], where the weight function $w(x_{1},x_{2})=\sin{(\pi x_{1})}\sin{(\pi x_{2})}$ on the unit square was used. As this weight is in $H^{1}_{0}(\Omega)\cap H^{2}(\Omega)$ almost first order convergence in time was observed, which we again can see for the Stokes optimal control problem. In the second example, we study an example where less regularity of the data is available, leading to a reduced order of convergence.
8.1 Solution Algorithm
Equation (6.5) can be written in matrix notation as
$\min_{{\bf{q}}_{\sigma}\in\mathbb{R}^{N}}\frac{1}{2}(S_{kh}M_{kh}^{uq}{\bf{q}}_{\sigma}-{\bf{u}}_{d,kh})^{T}M_{kh}^{uu}(S_{kh}M_{kh}^{uq}{\bf{q}}_{\sigma}-{\bf{u}}_{d,kh})+\frac{1}{2}\alpha{\bf{q}}_{\sigma}^{T}M_{kh}^{qq}{\bf{q}}_{\sigma}\qquad\text{s.t.}\quad W_{kh}S_{kh}{M_{kh}^{uq}}{\bf{q}}_{\sigma}\leq\beta,$ (8.1)
where $M_{kh}^{uu}$ is the mass matrix of $X^{0}_{k}(\mathbf{V}_{h})$, $M_{kh}^{qq}$ is the mass matrix of $\mathbf{Q}_{0}$, $M_{kh}^{uq}$ is the mass matrix encoding inner products of $X^{0}_{k}(\mathbf{V}_{h})$ and $\mathbf{Q}_{0}$ functions, ${\bf{u}}_{d,kh}$ is the $L^{2}$ projection of ${\bf{u}}_{d}$ onto $X^{0}_{k}(\mathbf{V}_{h})$ and $W_{kh}$ is the matrix that maps each ${\bf{u}}_{kh}$ to the vector $(G_{\bf{w}}({\bf{u}}_{kh})|_{I_{m}})_{m=1,..,M}$. Note that with a slight abuse of notation, we write $S_{kh}$ for the control to state mapping as well as the matrix representing this mapping. The optimality conditions for this problem read
$\alpha M_{kh}^{qq}{\bf{q}}_{\sigma}+(M_{kh}^{uq})^{T}S_{kh}^{T}M_{kh}^{uu}(S_{kh}M_{kh}^{uq}{\bf{q}}_{h}-{\bf{u}}_{kh,d})+(M_{kh}^{uq})^{T}S_{kh}^{T}W_{kh}^{T}\mu_{\sigma}$ $=0,$
$W_{kh}S_{kh}M_{kh}^{uq}{\bf{q}}_{\sigma}\leq\beta,\quad\mu_{\sigma}\geq 0,\quad\mu_{\sigma}^{T}(W_{kh}S_{kh}M_{kh}^{uq}{\bf{q}}_{\sigma}-\beta)$ $=0.$
We solve the above problem with a primal-dual-active-set strategy (PDAS), during which, for each iteration of the active set $\mathcal{A}_{n}\subset\{1,\dots,M\}$, we solve a symmetric saddle point system
$\begin{pmatrix}\alpha M_{kh}^{qq}+(M_{kh}^{uq})^{T}S_{kh}^{T}M_{kh}^{uu}S_{kh}M_{kh}^{uq}&(M_{kh}^{uq})^{T}S_{kh}^{T}W_{kh}^{T}O_{n}\\ O_{n}^{T}W_{kh}S_{kh}M_{kh}^{uq}&0\end{pmatrix}\begin{pmatrix}{\bf{q}}_{\sigma}^{n}\\ \mu_{\sigma}^{n}\end{pmatrix}=\begin{pmatrix}(M_{kh}^{uq})^{T}S_{kh}^{T}M_{kh}^{uu}{\bf{u}}_{kh,d}\\ \beta\cdot\mathds{1}_{|\mathcal{A}_{n}|}\end{pmatrix},$
where $O_{n}\in\mathbb{R}^{|\mathcal{A}_{n}|\times M}$ is the matrix satisfying $O_{n}\mu_{\sigma}=(\mu_{\sigma})_{\mathcal{A}_{n}}$, i.e., selects the active indices. We solve the linear system with MINRES, using the block diagonal preconditioner
$P_{kh}=\begin{pmatrix}\alpha M_{kh}^{qq}&0\\ 0&-\frac{1}{\alpha}W_{kh}S_{kh}M_{kh}^{uq}(M_{kh}^{qq})^{-1}(M_{kh}^{uq})^{T}S_{kh}^{T}W_{kh}^{T}\end{pmatrix}.$
Note that due to the choice of control discretization $M_{kh}^{qq}$ is a diagonal matrix, and if the partition of $I$ is uniform, and all mesh elements of $\Omega$ have the same volume, $M_{kh}^{qq}$ is a multiple of the identity matrix. Further note, that if the partition of $I$ is uniform, the matrix $S_{kh}^{T}W_{kh}^{T}$ has the structure
$\begin{pmatrix}{\bf{z}}_{M}&{\bf{z}}_{M-1}&...&{\bf{z}}_{2}&{\bf{z}}_{1}\\ 0&{\bf{z}}_{M}&...&{\bf{z}}_{3}&{\bf{z}}_{2}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&...&{\bf{z}}_{M}&{\bf{z}}_{M-1}\\ 0&0&...&0&{\bf{z}}_{M}\end{pmatrix}\in\mathbb{R}^{M\dim(\mathbf{V}_{h})\times M}$
where each vector ${\bf{z}}_{m}\in\mathbb{R}^{\dim(\mathbf{V}_{h})}$, $m=1,...,M$ corresponds to the degrees of freedom of ${\bf{z}}|_{I_{m}}$ for the solution ${\bf{z}}$ to
$B({\bf{v}}_{kh},{\bf{z}}_{kh})=\left({\bf{w}}\chi_{I_{M}},{\bf{v}}_{kh}\right)_{I\times\Omega}\quad\text{for all}\quad{\bf{v}}_{kh}\in X^{0}_{k}(\mathbf{V}_{h}).$
Hence, to assemble this matrix only one discrete adjoint problem has to be solved, and a decomposition of the preconditioning matrix $P_{kh}$ can be computed in advance and be reused in every iteration of the PDAS algorithm. The discrete solutions of the finite element problems were carried out in FEniCS Version 2019 [26], using the MINI Element in space.
8.2 Example 1
For this example, we consider the setting $\Omega=(0,1)^{2}$, $I=(0,1]$, choose the regularization parameter $\alpha=1$ and set the control constraints to ${\bf{q}}_{a}=(-\infty,-\infty)^{T}$, ${\bf{q}}_{b}=(+\infty,+\infty)^{T}$. We construct an analytic test case by considering the functions
$\varphi(t):=\begin{cases}{}48t^{2}-128t^{3}&t\in[0,1/4),\\ 1&t\in[1/4,3/4],\\ 48(1-t)^{2}-128(1-t)^{3}&t\in(3/4,1],\end{cases}\quad{\bf{y}}=\frac{64\sqrt{2}}{5\sqrt{7}}\begin{pmatrix}-\sin(\pi x_{1})^{4}\cos(\pi x_{2})\sin(\pi x_{2})^{3}\\ \cos(\pi x_{1})\sin(\pi x)^{3}\sin(\pi x_{2})^{4}\end{pmatrix}.$
Here, ${\bf{y}}$ has been constructed in such a way, that $\nabla\cdot{\bf{y}}=0$ and $\|{\bf{y}}\|_{L^{2}(\Omega)}=1$. It was obtained by considering the potential $\rho(x_{1},x_{2})=\left(\sin(\pi x_{1})\sin(\pi x_{2})\right)^{4}$, and normalizing the vector field $(\partial_{x_{2}}\rho(x_{1},x_{2}),-\partial_{x_{1}}\rho(x_{1},x_{2}))^{T}$. We choose $\bar{{\bf{u}}}=\varphi(t){\bf{y}}(x_{1},x_{2})$, ${\bf{w}}={\bf{y}}$ and $\beta\equiv 1$. It is then straightforward to verify, that $G_{\bf{w}}(\bar{{\bf{u}}})\leq\beta$ for all $t\in I$ and $G_{\bf{w}}(\bar{{\bf{u}}})=\beta$ if and only if $t\in[1/4,3/4]$. We thus choose the multiplier $\bar{\mu}=10^{3}\chi_{[1/4,3/4]}(t)$, which by construction satisfies $\bar{\mu}\geq 0$ and $\langle G_{\bf{w}}(\bar{\bf{u}})-\beta,\bar{\mu}\rangle=0$. We then proceed by choosing $\bar{p}=0$ and as a consequence set $\bar{\bf{q}}=\partial_{t}\bar{\bf{u}}-\Delta\bar{\bf{u}}$. We obtain $\bar{\bf{z}}=-\bar{\bf{q}}$ and with the choice $\bar{r}=0$ can fix ${\bf{u}}_{d}=\bar{\bf{u}}+\partial_{t}\bar{\bf{z}}+\Delta\bar{\bf{z}}+\bar{\mu}{\bf{w}}$, in such a way, that the constructed $(\bar{\bf{q}},\bar{\bf{u}})$ satisfy the first order optimality condition for this desired state. Note that $\vec{\psi}$ has been chosen sufficiently smooth at the boundary, such that $\Delta\vec{\psi}|_{\partial\Omega}=0$ and thus $\bar{\bf{q}}|_{\partial\Omega}=\bar{\bf{z}}|_{\partial\Omega}=0$. The calculation of the analytic solution was verified using the SageMath software [33]. We discretize this problem with a uniform triangulation of $\Omega$ and a uniform partition of $I$. To get more insight into the observed orders of convergence, for sequences of discretization levels $\{h_{l}\}$,$\{k_{l}\}$ we report the empirical orders of convergence determined by
$(\mathrm{EOC}_{h})_{l}$ $:=\dfrac{\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{k_{l}h_{l}}\|_{L^{2}(I;L^{2}(\Omega))})-\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{k_{l-1}h_{l-1}}\|_{L^{2}(I;L^{2}(\Omega))})}{\log(h_{l})-\log(h_{l-1})}$
$(\mathrm{EOC}_{k})_{l}$ $:=\dfrac{\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{k_{l}h_{l}}\|_{L^{2}(I;L^{2}(\Omega))})-\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{k_{l-1}h_{l-1}}\|_{L^{2}(I;L^{2}(\Omega))})}{\log(k_{l})-\log(k_{l-1})}$
Figure 1(a) displays convergence with respect to the spacial discretization parameter for fixed $k$. The theoretical convergence order of 1 can be observed. Figure 1(b) depicts the convergence results for the time discretization parameter for different values of $h$. Note that for this comparison, we have choosen discretization levels in time, such that the two boundary of the active set $[1/4,3/4]$ are midpoints of the discrete subintervals, in order to exclude superconvergence effects. Due to the structure of the discretization, the number of degrees of freedom, grows rather quickly, which is why very fine discretizations are expensive. Figure 1(b) shows, that if the spacial discretization parameter is chosen too large, the stagnation phase sets in rather early. If one only observes the coarse discretizations, the observed order of convergence is skewed, which led to an estimated order of convergence of 0.85 reported in [7]. As the derivation of the present example with an analytic reference solution requires the involved functions to have some precise regularity, we next propose an example with desired state that has low regularity in time. In this case a numerical reference solution is needed to measure the errors.
$k=10^{-3}$
$h$ err $\mathrm{EOC}_{h}$
$2^{-3}$ 36.954 -
$2^{-4}$ 16.899 1.12
$2^{-5}$ 8.0461 1.07
$2^{-6}$ 3.9657 1.02
((a)) Convergence with respect to $h$ for $k=10^{-3}$.
$h=2^{-6}$ $h=2^{-7}$ $h=2^{-8}$
$k$ err $\mathrm{EOC}_{k}$ err $\mathrm{EOC}_{k}$ err $\mathrm{EOC}_{k}$
$6^{-1}$ 15.373 - 14.977 - 14.878 -
$18^{-1}$ 6.5178 0.78 5.5318 0.90 5.2592 0.94
$54^{-1}$ 4.3312 0.37 2.6310 0.67 1.9980 0.88
$162^{-1}$ 4.0066 0.07 2.0579 0.22 1.1443 0.50
((b)) Convergence with respect to $k$ for different values of $h$.
Figure 1: Numerical observation of the error $\|\bar{\bf{q}}-\bar{\bf{q}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}$ for Example 1.
8.3 Example 2
In this example we consider an optimal control problem where we specify rough data ${\bf{u}}_{d}$, and ${\bf{w}}$, while keeping the domain $\Omega=(0,1)^{2}$, $I=(0,1]$ and ${\bf{q}}_{a}=(-\infty,-\infty)^{T}$, ${\bf{q}}_{b}=(+\infty,+\infty)^{T}$. We specify
${\bf{u}}_{d}=5\cdot 10^{4}\varphi(t)\begin{pmatrix}\sin(\pi x_{2})\cos(\pi x_{2})\sin(\pi x_{1})^{2}\\ -\sin(\pi x_{1})\cos(\pi x_{1})\sin(\pi x_{2})^{2}\end{pmatrix}\quad\text{where}\quad\varphi(t)=\begin{cases}{}\sqrt{t-\frac{1}{5}}\cdot(\frac{2}{5}-t)&\text{if }t\in[\frac{1}{5},\frac{2}{5}],\\ -\sqrt{t-\frac{3}{5}}\cdot(\frac{4}{5}-t)&\text{if }t\in[\frac{3}{5},\frac{4}{5}],\\ 0&\text{else.}\end{cases}$
The weight in the state constraint is given by
${\bf{w}}=\chi_{\{(x_{1}-0.5)^{2}+(x_{2}-0.5)^{2}\leq 0.125\}}(x_{1},x_{2})\cdot\begin{pmatrix}x_{2}-0.5\\ -x_{1}+0.5\end{pmatrix}$
and the scalar constraint is given by $\beta\equiv 1$. Moreover, we consider the regularization parameter $\alpha=10^{-4}$. As in this case, no analytical optimal solution is known, we estimate the errors using a numerical reference solution on a fine grid. To this end, we discretize the problem on 960 time intervals and 128 subdivisions of $\Omega$ in each direction. Note that due to this evaluation of the errors, we expect a faster convergence than theoretically derived, as on the finest discretization level, the error would equal to 00. In Figure 3(a) we can again observe order 1 convergence with respect to $h$ for fixed time discretization. In Figure 3(b) we observe that the convergence with respect to $k$ exhibits a rate of about $0.6$, which is much closer to the analytically derived $0.5$ than the rate observed for the smooth example 1.
((a)) Time function $\varphi(t)$.
((b)) Weight function ${\bf{w}}$.
((c)) Satisfaction of the state constraint.
Figure 2: Data for Example 2.
$h$ err $\mathrm{EOC}_{h}$
$2^{-2}$ 417.33 -
$2^{-3}$ 205.87 1.01
$2^{-4}$ 96.306 1.09
$2^{-5}$ 42.372 1.18
((a)) Convergence with respect to $h$ for $k=960^{-1}$.
$k$ err $\mathrm{EOC}_{k}$
$5^{-1}$ 399.06 -
$10^{-1}$ 265.39 0.58
$20^{-1}$ 175.97 0.59
$40^{-1}$ 118.80 0.56
$80^{-1}$ 74.697 0.66
((b)) Convergence with respect to $k$ for $h=2^{-7}$.
Figure 3: Numerical observation of the error $\|\bar{\bf{q}}-\bar{\bf{q}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}$ for Example 2.

### v2

@@
 The discretization scheme for the Stokes equations consists of inf-sup stable finite elements in space and a discontinuous Galerkin method in time, for which we have recently established best approximation type error estimates.
-Using these error estimates, for the discrete control problem we establish error estimates and as a by-product we show an improved regularity for the optimal control.
+Using these error estimates, for the discrete control problem we derive error estimates and as a by-product we show an improved regularity for the optimal control.
 We complement our theoretical analysis with numerical results.
 keywords:
-keyword1; Keyword2; Keyword3; Keyword4.
+Stokes equations, instationary, optimal control, state constraints, error estimates
 1 Introduction
@@
 Here we assume that $\Omega\subset\mathbb{R}^{d}$, $d=2,3$, is a convex polygonal or polyhedral domain and $I=(0,T]$ is a bounded time interval.
+In the objective function, ${\bf{u}}_{d}\in L^{2}(I\times\Omega)^{d}$ represents the desired state and $\alpha>0$ is the regularization parameter.
 The control constraints are given by the constant vectors ${\bf{q}}_{a},{\bf{q}}_{b}\in{(\mathbb{R}\cup\{\pm\infty\})^{d}}$ and satisfy ${\bf{q}}_{a}<{\bf{q}}_{b}$.
-The constant scalar $\beta$ satisfies $\beta>0$, ${\bf w}(x)$ is a given function in $L^{2}(\Omega)^{d}$ and $\alpha>0$ is some regularization parameter.
+In the state constraint, the constant scalar $\beta$ satisfies $\beta>0$ and ${\bf w}(x)$ is a given function in $L^{2}(\Omega)^{d}$.
 Note that for ease of presentation, we consider an optimal control problem with homogeneous initial data ${\bf{u}}(0)={\bf{0}}$, while all results also extend to the inhomogeneous case ${\bf{u}}(0)={\bf{u}}_{0}$.
 The main result of this paper states, that the error between the optimal control $\bar{{\bf{q}}}$ for the continuous problem and the optimal solution $\bar{{\bf{q}}}_{\sigma}$ of the discretized problem satisfies
-$\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{\sigma}\rVert_{L^{2}(I;L^{2}(\Omega))}\leq\frac{C}{\sqrt{\alpha}}\,\ln{\frac{T}{k}}(k^{\frac{1}{2}}+h),$
+$\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{\sigma}\rVert_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq\frac{C}{\sqrt{\alpha}}\,\ln{\frac{T}{k}}(k^{\frac{1}{2}}+h),$
 […]
@@
 We denote for a Banach space $X$ and $1\leq p\leq\infty$, by $L^{p}(I;X)$ the Bochner space of $X$-valued functions over $I$, whose $X$-norm is $p$-integrable w.r.t time.
-If $X$ is reflexive and $1\leq p<\infty$, there holds the following isomorphism, see [20, Corollary 1.3.22]
+If $X$ is reflexive and $1\leq p<\infty$, there holds the following isomorphism, see [23, Corollary 1.3.22]
 $L^{p}(I;X)^{*}\cong L^{q}(I;X^{*}),\qquad\text{where }\frac{1}{p}+\frac{1}{q}=1.$ (2.1)
@@
 where by ${\bf{v}}\cdot\mathbf{n}$ we denote the generalized normal trace. We denote by $\mathbf{V}^{*}$ the topological dual space of $\mathbf{V}$, and define
-$\mathbf{W}$ $:=L^{2}(I;\mathbf{V})\cap H^{1}(I;\mathbf{V}^{*})\hookrightarrow C(\bar{I};{\mathbf{V}_{0}}),$ (2.2)
+$\mathbf{W}$ $:=L^{2}(I;\mathbf{V})\cap H^{1}(I;\mathbf{V}^{*})\hookrightarrow C(\bar{I};{\mathbf{V}_{0}}),$
 $\mathbf{X}$ :=$:=\{{\bf{v}}\in L^{2}(I;\mathbf{V}):\partial_{t}{\bf{v}}\in L^{2}(I;\mathbf{V}^{*})+L^{1}(I;\mathbf{V}_{0})\text{ and }{\bf{v}}(0)={\bf{0}}\},$
@@
 as it holds $\|A^{\frac{1}{2}}{\bf{u}}\|_{L^{2}(\Omega)}^{2}=\left(A{\bf{u}},{\bf{u}}\right)_{\Omega}=\left(\nabla{\bf{u}},\nabla{\bf{u}}\right)_{\Omega}=\|\nabla{\bf{u}}\|_{L^{2}(\Omega)}^{2}$.
-For the proof of $D(A^{\frac{1}{2}})=\mathbf{V}$, see [31, Ch.
+For the proof of $D(A^{\frac{1}{2}})=\mathbf{V}$, see [35, Ch.
 III, Lemma 2.2.1].
@@
 If ${\bf{q}}\in L^{2}(I;{L^{2}(\Omega)^{d}})$, then ${\bf{u}}\in L^{2}(I;{\mathbf{V}_{2}})\cap H^{1}(I;{\mathbf{V}_{0}})\hookrightarrow C(\bar{I};\mathbf{V})$, and there holds
-$\|{\bf{u}}\|_{L^{2}(I;H^{2}(\Omega))}+\|{\bf{u}}\|_{L^{\infty}(I;\mathbf{V})}+\|\partial_{t}{\bf{u}}\|_{L^{2}(I;L^{2}(\Omega))}\leq C\|{\bf{q}}\|_{L^{2}(I;L^{2}(\Omega))}.$
+$\|{\bf{u}}\|_{L^{2}(I;H^{2}(\Omega))}+\|{\bf{u}}\|_{L^{\infty}(I;\mathbf{V})}+\|\partial_{t}{\bf{u}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq C\|{\bf{q}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}.$
 Proof.
-The first part of this theorem is proven in [32, Chapter III, Theorem 1.1] and the remark on page 179 therein.
-The $H^{2}$ regularity part can be shown as in [32, Chapter III, Proposition 1.2] using the $H^{2}$ regularity result for the stationary Stokes equations in convex polygonal/polyhedral domains, proven in [9, Theorem 5.5, Theorem 6.3] see also [23, Theorem 2].
+The first part of this theorem is proven in [36, Chapter III, Theorem 1.1] and the remark on page 179 therein.
+The $H^{2}$ regularity part can be shown as in [36, Chapter III, Proposition 1.2] using the $H^{2}$ regularity result for the stationary Stokes equations in convex polygonal/polyhedral domains, proven in [10, Theorem 5.5, Theorem 6.3] see also [26, Theorem 2].
 ∎
@@
 […]
-see [3, Proposition 2.6], holding on fairly general domains, e.g.
+see [4, Proposition 2.6], holding on fairly general domains, e.g.
 Lipschitz domains.
@@
 Most often maximal parabolic regularity is treated in $L^{2}$ or $L^{q}$ setting in space, but it can also be extended to settings of weaker spacial regularity.
-The property (2.3) combined with [2, Lemma 11.4] yields that maximal parabolic regularity also holds in $\mathbf{V}^{*}$, i.e.
+The property (2.2) combined with [3, Lemma 11.4] yields that maximal parabolic regularity also holds in $\mathbf{V}^{*}$, i.e.
 […]
@@
 The regularity of the pressure depends on the regularity of the right hand side and the velocity component of the solution.
-There holds the following result, see [3, Theorem 2.10, Corollary 2.11].
+There holds the following result, see [4, Theorem 2.10, Corollary 2.11].
 Proposition 2.2.
@@
 Note, that the regularity $\partial_{t}{\bf{u}}\in L^{2}(I;\mathbf{V}^{*})$ can only be obtained by bootstrapping if the right hand side is in $L^{2}(I;\mathbf{V}^{*})$.
-The third claim is obtained by using the maximal parabolic regularity and $H^{2}$ regularity of the stationary Stokes problem on convex polygonal/polyhedral domains, see [9, Theorem 5.5, Theorem 6.3].
+The third claim is obtained by using the maximal parabolic regularity and $H^{2}$ regularity of the stationary Stokes problem on convex polygonal/polyhedral domains, see [10, Theorem 5.5, Theorem 6.3].
 ∎
@@
 The results presented in this work however also hold true in the inhomogeneous initial data case.
-In this case, the assumption $\beta>0$ then has to be generalized to $\beta>\left({\bf{u}}_{0},{\bf{w}}\right)_{\Omega}$.
+In this case, the assumption $\beta>0$ , on the parameter $\beta$ of the state constraint, then has to be generalized to $\beta>\left({\bf{u}}_{0},{\bf{w}}\right)_{\Omega}$.
 Let us next give a characterization of the adjoint operator $S^{*}$.
-Theorem 2.1, together with [5, Proposition V.1.3], yields that for any ${\bf{q}}\in{\mathbf{Y}}$, there exists a unique solution ${\bf{u}}\in\mathbf{X}$ such that
+Theorem 2.1, together with [6, Proposition V.1.3], yields that for any ${\bf{q}}\in{\mathbf{Y}}$, there exists a unique solution ${\bf{u}}\in\mathbf{X}$ such that
 $a({\bf{u}},{\bf{v}}):=\left\langle\partial_{t}{\bf{u}},{\bf{v}}\right\rangle_{I\times\Omega}+\left(\nabla{\bf{u}},\nabla{\bf{v}}\right)_{I\times\Omega}=\left\langle{\bf{q}},{\bf{v}}\right\rangle_{I\times\Omega}\qquad\forall{\bf{v}}\in{\mathbf{Y}^{*}},$ (3.1)
@@
 and let $\{{\bf{q}}_{n}\}$ and $\{{\bf{u}}_{n}:={\bf{u}}({\bf{q}}_{n})\}$ denote sequences of feasible controls with associated states, such that $J({\bf{q}}_{n},{\bf{u}}_{n})\to\mathcal{J}$ as $n\to\infty$.
-As it holds $J({\bf{q}}_{n},{\bf{u}}_{n})\leq{\mathcal{J}+1}$ for large enough $n$, there holds a bound $\|{\bf{q}}_{n}\|_{L^{2}(I;L^{2}(\Omega))}\leq C$ for all $n$.
+As it holds $J({\bf{q}}_{n},{\bf{u}}_{n})\leq{\mathcal{J}+1}$ for large enough $n$, there holds a bound $\|{\bf{q}}_{n}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq C$ for all $n$.
 From Theorem 3.1, we obtain $\|{\bf{u}}_{n}\|_{L^{2}(I;H^{2}(\Omega))}+\|{\bf{u}}_{n}\|_{H^{1}(I;L^{2}(\Omega))}\leq C$.
@@
 These allow us to pass to the limit in the weak form of the state equation, showing that $\hat{\bf{u}}={\bf{u}}(\hat{\bf{q}})$.
-Furthermore, as $L^{2}(I;H^{2}(\Omega)^{d})\cap H^{1}(I;L^{2}(\Omega)^{d})\hookrightarrow C({\bar{I}};H^{s}(\Omega)^{d})$ compactly, for $s<1$, see [30, Corollary 8], by taking another subsequence, we obtain ${\bf{u}}_{n}\to\hat{\bf{u}}$ in $C({\bar{I}};L^{2}(\Omega)^{d})$.
+Furthermore, as $L^{2}(I;H^{2}(\Omega)^{d})\cap H^{1}(I;L^{2}(\Omega)^{d})\hookrightarrow C({\bar{I}};H^{s}(\Omega)^{d})$ compactly, for $s<1$, see [34, Corollary 8], by taking another subsequence, we obtain ${\bf{u}}_{n}\to\hat{\bf{u}}$ in $C({\bar{I}};L^{2}(\Omega)^{d})$.
 As $G_{\bf{w}}({\bf{u}}_{n})\leq\beta$ for all $t\in I$, this shows $G_{\bf{w}}(\hat{\bf{u}})\leq\beta$ for all $t\in I$.
@@
 By the generalized KKT condition (cf.
-[6, Theorem 5.2]) and the Slater condition from Assumption 1, i.e., the existence of $\tilde{{\bf{q}}}\in\mathbf{Q}_{ad}$ such that $\mathcal{G}(\tilde{{\bf{q}}})\in\operatorname{int}\mathcal{K}$, the optimality of $\bar{{\bf{q}}}$ is equivalent to the existence of a Lagrange multiplier $\bar{\mu}\in C(\bar{I})^{*}$ and the adjoint state
+[7, Theorem 5.2]) and the Slater condition from Assumption 1, i.e., the existence of $\tilde{{\bf{q}}}\in\mathbf{Q}_{ad}$ such that $\mathcal{G}(\tilde{{\bf{q}}})\in\operatorname{int}\mathcal{K}$, the optimality of $\bar{{\bf{q}}}$ is equivalent to the existence of a Lagrange multiplier $\bar{\mu}\in C(\bar{I})^{*}$ and the adjoint state
 $\bar{{\bf{z}}}=S^{\prime}(\bar{{\bf{q}}})^{*}(S(\bar{{\bf{q}}})-{\bf{u}}_{d})+\mathcal{G}^{\prime}(\bar{{\bf{q}}})^{*}\bar{\mu}$ (3.7)
@@
 Using [1, Theorem 4.31], there exists a normalized function of bounded variation $\tilde{\mu}\in\mathrm{NBV}(\bar{I})$, such that the application of $\mu\in C(\bar{I})^{*}$ to any $\xi\in C(\bar{I})$ can be expressed as a Riemann-Stieltjes integral
-$\left\langle\xi,\mu\right\rangle_{I\times\Omega}=\int_{I}\xi(t)\ \mathrm{d}\tilde{\mu}(t).$
+${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left\langle\xi,\mu\right\rangle_{I}}=\int_{I}\xi(t)\ \mathrm{d}\tilde{\mu}(t).$
 Slightly modifying [1, Definition 1.2], we can normalize $\tilde{\mu}$ such that it is of bounded variation, right continuous and satisfies $\tilde{\mu}(T)=0$.
 If $\xi\in C^{1}(\bar{I})$, due to [1, Proposition 4.24, Theorem 4.17], there holds
-$\left\langle\xi,\mu\right\rangle_{I\times\Omega}=-\int_{I}\tilde{\mu}(t)\ \mathrm{d}\xi(t)+\tilde{\mu}(T)\xi(T)-\tilde{\mu}(0)\xi(0)=-\int_{I}\tilde{\mu}(t)\partial_{t}\xi(t)\ \mathrm{d}t-\tilde{\mu}(0)\xi(0)$
+${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left\langle\xi,\mu\right\rangle_{I}}=-\int_{I}\tilde{\mu}(t)\ \mathrm{d}\xi(t)+\tilde{\mu}(T)\xi(T)-\tilde{\mu}(0)\xi(0)=-\int_{I}\tilde{\mu}(t)\partial_{t}\xi(t)\ \mathrm{d}t-\tilde{\mu}(0)\xi(0)$
 where the last integral can be understood equivalently in the Riemann or Lebesgue sense.
@@
 $[\varphi](t)=\varphi(t+)-\varphi(t-)$.
-Since functions in $\mathrm{BV}(I;X)$ for any Banach space $X$, possess well-defined onesided limits, see [19, Propositions 2.1 & 2.2], we obtain from (3.10) that ${\bf{z}}_{\alpha}(t\pm)\in\mathbf{V}_{0}$ is well defined for any $t\in I$.
+Since functions in $\mathrm{BV}(I;X)$ for any Banach space $X$, possess well-defined onesided limits, see [20, Propositions 2.1 & 2.2], we obtain from (3.10) that ${\bf{z}}_{\alpha}(t\pm)\in\mathbf{V}_{0}$ is well defined for any $t\in I$.
 Hence, by continuity of $P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}\colon L^{2}(\Omega)^{d}\to L^{2}(\Omega)^{d}$ we also have $[\bar{\bf{q}}]=[P_{[{\bf{q}}_{a},{\bf{q}}_{b}]}({\bf{z}}_{\alpha})]\in L^{2}(\Omega)^{d}$ for all $t\in I$.
@@
 As ${\bf{w}}\in\mathbf{V}$, due to (2.3), it holds
-$\left(\bar{\bf{q}},{\bf{w}}\right)_{\Omega}+\left(\nabla\bar{\bf{u}},\nabla{\bf{w}}\right)_{\Omega}=\left(\partial_{t}\bar{\bf{u}},{\bf{w}}\right)_{\Omega}=\partial_{t}G_{\bf{w}}(\bar{\bf{u}})\quad\text{a.e.
+$\left(\bar{\bf{q}},{\bf{w}}\right)_{\Omega}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}-}\left(\nabla\bar{\bf{u}},\nabla{\bf{w}}\right)_{\Omega}=\left(\partial_{t}\bar{\bf{u}},{\bf{w}}\right)_{\Omega}=\partial_{t}G_{\bf{w}}(\bar{\bf{u}})\quad\text{a.e.
 in }I.$
@@
 $\hat{\bf{z}}\in W^{1,s}(I;\mathbf{V}_{0})\cap L^{s}(I;\mathbf{V}_{2})\hookrightarrow C({\bar{I}};C(\bar{\Omega})),$
-where the last embedding holds due to [30, Corollary 8].
+where the last embedding holds due to [34, Corollary 8].
 As further ${\bf{w}}\in\mathbf{V}_{2}\hookrightarrow C(\bar{\Omega})$, for any $t^{*}\in I$ it thus holds $\bar{\bf{z}}(t^{*}+),\bar{\bf{z}}(t^{*}-)\in C(\bar{\Omega})$.
@@
 $\|q-r_{h}q\|_{L^{2}(\Omega)}$ $\leq ch\|\nabla q\|_{L^{2}(\Omega)}$ $\forall q\in H^{1}(\Omega).$
-This assumption is valid, for example, for Taylor-Hood and MINI finite elements on shape regular meshes, see [3, Assumption 7.2].
-We introduce the vectorial discrete Laplace operator $-\Delta_{h}\colon\mathbf{X}_{h}\rightarrow\mathbf{X}_{h}$ by
-$(-\Delta_{h}{\bf{v}}_{h},\chi)=(\nabla{\bf{v}}_{h},\nabla\chi),\quad\forall{\bf{v}}_{h},\chi\in\mathbf{X}_{h}.$ (4.2)
-We also define the space of discretely divergence-free vector fields $\mathbf{V}_{h}$ as
+This assumption is valid, for example, for Taylor-Hood and MINI finite elements on shape regular meshes, see [4, Assumption 7.2].
+We define the space of discretely divergence-free vector fields $\mathbf{V}_{h}$ as
 $\mathbf{V}_{h}=\Set{{\bf{v}}_{h}\in\mathbf{X}_{h}\colon(\nabla\cdot{\bf{v}}_{h},q_{h})=0\quad\forall q_{h}\in M_{h}}.$ (4.2)
@@
 $\left(\nabla({\bf{u}}-{\bf{u}}_{h}),\nabla{\bf{v}}_{h}\right)_{\Omega}-\left(\nabla\cdot{\bf{v}}_{h},(p-p_{h})\right)_{\Omega}+\left(\nabla\cdot({\bf{u}}-{\bf{u}}_{h}),q_{h}\right)_{\Omega}=0\quad\forall({\bf{v}}_{h},q_{h})\in\mathbf{X}_{h}\times M_{h}.$ (4.5)
-Note that, if ${\bf{u}}\in\mathbf{V}$, then it holds $R_{h}^{S}({\bf{u}},p)\in\mathbf{V}_{h}$. Further, the Stokes Ritz projection satisfies the following stability, see [4, Theorem 5.2.1].
+Note that, if ${\bf{u}}\in\mathbf{V}$, then it holds $R_{h}^{S}({\bf{u}},p)\in\mathbf{V}_{h}$. Further, the Stokes Ritz projection satisfies the following stability, see [5, Theorem 5.2.1].
 $\|\nabla R_{h}^{S}({\bf{u}},p)\|_{L^{2}(\Omega)}+\|R_{h}^{S,p}({\bf{u}},p)\|_{L^{2}(\Omega)}\leq C(\|\nabla{\bf{u}}\|_{L^{2}(\Omega)}+\|p\|_{L^{2}(\Omega)}).$
-Let us recall the following error estimates for the stationary discrete Stokes problem: See [4, Theorem 5.25], [22, Theorems 4.21, 4.25, 4.28], [15, Theorems 53.17 & 53.19] or [17, Chapter II, Theorems 1.8 & 1.9].
+Let us recall the following error estimates for the stationary discrete Stokes problem: See [5, Theorem 5.25], [25, Theorems 4.21, 4.25, 4.28], [16, Theorems 53.17 & 53.19] or [18, Chapter II, Theorems 1.8 & 1.9].
 Theorem 4.1.
@@
 $\|{\bf{u}}-{\bf{u}}_{h}\|_{L^{2}(\Omega)}\leq Ch(\|{\bf{u}}-{\bf{u}}_{h}\|_{H^{1}(\Omega)}+\|p-p_{h}\|_{L^{2}(\Omega)}).$
-Note, that while [4, Theorem 5.5.6] contains formal results, on how to derive error estimates also for $p$ in a weaker norm, e.g.
+Note, that while [5, Theorem 5.5.6] contains formal results, on how to derive error estimates also for $p$ in a weaker norm, e.g.
 $H^{-1}(\Omega)$, the argument requires $H^{2}$ regularity results for the compressible Stokes equations.
@@
 Let ${\bf{q}}\in L^{2}(I;H^{-1}(\Omega)^{d})+L^{1}(I;L^{2}(\Omega)^{d})$. Then there exists a unique solution ${\bf{u}}_{kh}\in X^{0}_{k}(\mathbf{V}_{h})$ of (4.10), satsifying
-$\|{\bf{u}}_{kh}\|_{L^{2}(I;H^{1}(\Omega))}+\|{\bf{u}}_{kh}\|_{L^{\infty}(I;L^{2}(\Omega))}\leq C\|{\bf{q}}\|_{L^{2}(I;H^{-1}(\Omega))+L^{1}(I;L^{2}(\Omega))}.$
+$\|{\bf{u}}_{kh}\|_{L^{2}(I;H^{1}(\Omega))}+\|{\bf{u}}_{kh}\|_{L^{\infty}(I;L^{2}(\Omega))}\leq C\|{\bf{q}}\|_{L^{2}(I;H^{-1}(\Omega))+L^{1}(I;L^{2}(\Omega))},$
+where the constant $C>0$ is independent of $k,h$.
 Proof.
@@
 This is a standard argument, making use of a discrete Gronwall Lemma, see e.g.
-[34, Theorem 4.13] for a proof focusing especially on the $L^{1}(I;L^{2}(\Omega)^{d})$ right hand side case.
+[38, Theorem 4.13] for a proof focusing especially on the $L^{1}(I;L^{2}(\Omega)^{d})$ right hand side case.
 ∎
@@
 In our recent paper [4], we have established a best approximation type error estimate for the Stokes problem in the $L^{\infty}(I;L^{2}(\Omega)^{d})$ norm.
-From this more general result, we obtain in the case of homogeneous initial data the following result, see [3, Corollary 6.4].
+From this more general result, we obtain in the case of homogeneous initial data the following result, see [4, Corollary 6.4].
 Theorem 4.4.
@@
 where $\ell_{k}=\ln{\frac{T}{k}}$ and $R^{S}_{h}({\bf{u}},p)$ is the stationary finite element Stokes projection introduced in (4.5).
-Using the error estimates for the stationary Stokes Ritz projection of Theorem 4.1, in [3, Theorem 7.4] the following estimate in terms of explicit orders of convergence was shown.
+Using the error estimates for the stationary Stokes Ritz projection of Theorem 4.1, in [4, Theorem 7.4] the following estimate in terms of explicit orders of convergence was shown.
 Corollary 4.5.
@@
 However, using the energy and duality arguments it is possible to show the corresponding results log-free and with less regularity assumptions on the data (cf.
-[24, Theorem 11 & Theorem 13]).
+[28, Theorem 11 & Theorem 13]).
 Theorem 4.6.
@@
 If in addition to assumptions of Theorem 4.6, the domain $\Omega$ is convex, then there exists a constant $C$ independent of $k$ and $h$ such that
-$\|{\bf{u}}-{\bf{u}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}\leq C(h^{2}+k)\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{q}}}\|_{L^{2}(I;L^{2}(\Omega))}.$
+$\|{\bf{u}}-{\bf{u}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq C(h^{2}+k)\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{q}}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}.$
 5 Variational Discretization of the Optimal Control Problem
 In this section we consider the optimal control problem subject to the fully discretized Stokes equations.
-We consider a variational discretization for the controls, i.e., do not fix a finite dimensional approximation of the control space yet.
-The problem reads
-Minimize $\text{Minimize }\;J({\bf{q}}_{kh},{\bf{u}}_{kh})=\frac{1}{2}\lVert{\bf{u}}_{kh}-{\bf{u}}_{d}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\frac{\alpha}{2}\lVert{\bf{q}}_{kh}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}$ (5.1a)
+We consider a variational discretization for the controls, i.e., do not fix a finite dimensional approximation of the control space yet, cf., [15, 21] The problem reads
+Minimize $\text{Minimize }\;J({\bf{q}}_{kh},{\bf{u}}_{kh})=\frac{1}{2}\lVert{\bf{u}}_{kh}-{\bf{u}}_{d}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\frac{\alpha}{2}\lVert{\bf{q}}_{kh}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$ (5.1a)
 over all ${\bf{q}}_{kh}\in\mathbf{Q}_{ad}$, ${\bf{u}}_{kh}\in X^{0}_{k}(\mathbf{V}_{h})$, subject to
@@
 Discrete state constraint and complementarity conditions
-${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}G_{\bf{w}}(\bar{\bf{u}}_{kh})|_{I_{m}}\leq\beta,\ m=1,...,M},\quad\bar{\mu}_{kh}\geq 0\quad\text{and}\quad\langle\bar{\mu}_{kh},b-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}G_{\bf{w}}(\bar{\bf{u}}_{kh})}\rangle=0;$ (5.5b)
+${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}G_{\bf{w}}(\bar{\bf{u}}_{kh})|_{I_{m}}\leq\beta,\ m=1,...,M},\quad\bar{\mu}_{kh}\geq 0\quad\text{and}\quad\langle\bar{\mu}_{kh},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}G_{\bf{w}}(\bar{\bf{u}}_{kh})}\rangle=0;$ (5.5b)
 Discrete adjoint equation
@@
 Under 3 and for $h$ small enough, there exists $C>0$ independent of $k$, $h$, such that the optimal control $\bar{{\bf{q}}}_{kh}\in\mathbf{Q}_{ad}$ of the variationally discretized problem (5.1a) - (5.1c), together with its corresponding state $\bar{{\bf{u}}}_{kh}\in X^{0}_{k}(\mathbf{V}_{h})$ and corresponding multiplier $\bar{\mu}_{kh}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{1}(\bar{I})}$ satisfy the bound
-$\|\bar{\bf{q}}_{kh}\|_{L^{\infty}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{\mu}_{kh}\|_{L^{1}(\bar{I})}\leq C.$
+$\|\bar{\bf{q}}_{kh}\|_{L^{\infty}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\|\bar{\mu}_{kh}\|_{L^{1}(\bar{I})}\leq C.$
 Proof.
 Due to the feasibility of $\tilde{\bf{q}}$, shown in Lemma 5.2, it holds
-$J(\bar{{\bf{q}}}_{kh},\bar{{\bf{u}}}_{kh})\leq J(\tilde{{\bf{q}}},{\bf{u}}_{kh}(\tilde{{\bf{q}}}))$ $=\frac{1}{2}\lVert{\bf{u}}_{kh}(\tilde{{\bf{q}}})-{\bf{u}}_{d}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\frac{\alpha}{2}\lVert\tilde{{\bf{q}}}\rVert^{2}_{L^{2}(I)}\leq C(T,\tilde{{\bf{q}}}),$
+$J(\bar{{\bf{q}}}_{kh},\bar{{\bf{u}}}_{kh})\leq J(\tilde{{\bf{q}}},{\bf{u}}_{kh}(\tilde{{\bf{q}}}))$ $=\frac{1}{2}\lVert{\bf{u}}_{kh}(\tilde{{\bf{q}}})-{\bf{u}}_{d}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\frac{\alpha}{2}\lVert\tilde{{\bf{q}}}\rVert^{2}_{L^{2}(I)}\leq C(T,\tilde{{\bf{q}}}),$
 where due to Theorem 4.2, this bound is independent of $k$ and $h$. This results in
-$\|\bar{{\bf{q}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}\leq C.$ (5.7)
+$\|\bar{{\bf{q}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\|\bar{{\bf{u}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq C.$ (5.7)
 Let us define ${\bf{p}}=\frac{1}{2}\bar{{\bf{q}}}+\frac{1}{2}\tilde{{\bf{q}}}$.
@@
 Inserting this into (5.9), yields together with $\bar{\mu}_{kh}\geq 0$ and the complementarity conditions (5.5b):
-$0\leq C+\langle\bar{\mu}_{kh},b-G_{\bf{w}}(\bar{{\bf{u}}}_{kh})\rangle-\frac{\delta}{4}\langle\bar{\mu}_{kh},1\rangle=C-\frac{\delta}{4}\langle\bar{\mu}_{kh},1\rangle.$
+$0\leq C+\langle\bar{\mu}_{kh},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta}-G_{\bf{w}}(\bar{{\bf{u}}}_{kh})\rangle-\frac{\delta}{4}\langle\bar{\mu}_{kh},1\rangle=C-\frac{\delta}{4}\langle\bar{\mu}_{kh},1\rangle.$
 Thus, again using $\bar{\mu}_{kh}\geq 0$ and Remark 5.5, results in
@@
 Then there exists a constant $C>0$, such that it holds
-$\sqrt{\alpha}\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{kh}\rVert_{L^{2}(I;L^{2}(\Omega))}+\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{kh}\rVert_{L^{2}(I;L^{2}(\Omega))}\leq C\,\ell_{k}(k^{\frac{1}{2}}+h),\quad\ell_{k}=\ln{\frac{T}{k}}.$
+$\sqrt{\alpha}\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{kh}\rVert_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{kh}\rVert_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq C\,\ell_{k}(k^{\frac{1}{2}}+h),\quad\ell_{k}=\ln{\frac{T}{k}}.$
 Proof.
@@
 Adding these two inequalities results in
-$\alpha\|\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}\|^{2}_{L^{2}(I;L^{2}(\Omega))}\leq(\bar{{\bf{z}}},\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}})_{I\times\Omega}+(\bar{{\bf{z}}}_{kh},\bar{{\bf{q}}}-\bar{{\bf{q}}}_{kh})_{I\times\Omega}:=I_{1}+I_{2}.$ (5.12)
+$\alpha\|\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}\|^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq(\bar{{\bf{z}}},\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}})_{I\times\Omega}+(\bar{{\bf{z}}}_{kh},\bar{{\bf{q}}}-\bar{{\bf{q}}}_{kh})_{I\times\Omega}:=I_{1}+I_{2}.$ (5.12)
 We estimate the two terms separately.
@@
 $+\|P_{\beta}G_{\bf w}(\bar{{\bf{u}}}_{kh})-P_{\beta}G_{\bf w}({\bf{u}}(\bar{{\bf{q}}}_{kh})\|_{L^{\infty}(I)}\|\bar{\mu}\|_{C(\bar{I})^{*}}$
-$\leq$ $2\|{\bf w}\|_{L^{2}(\Omega)}\|\bar{{\bf{u}}}_{kh}(\bar{{\bf{q}}}_{kh})-{\bf{u}}(\bar{{\bf{q}}}_{kh})\|_{L^{\infty}(I;L^{2}(\Omega))}\|\bar{\mu}\|_{C(\bar{I})^{*}}$
+$\leq$ $2\|{\bf w}\|_{L^{2}(\Omega)}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{u}}_{kh}}(\bar{{\bf{q}}}_{kh})-{\bf{u}}(\bar{{\bf{q}}}_{kh})\|_{L^{\infty}(I;L^{2}(\Omega))}\|\bar{\mu}\|_{C(\bar{I})^{*}}$
 $\leq$ $C\ell_{k}^{2}(h^{2}+k)\|\bar{{\bf{q}}}_{kh}\|_{L^{\infty}(I;L^{2}(\Omega))}\|{\bf w}\|_{L^{2}(\Omega)}\|\bar{\mu}\|_{C(\bar{I})^{*}}$
@@
 $+\|P_{\beta}G_{\bf w}(\bar{{\bf{u}}})-P_{\beta}G_{\bf w}({\bf{u}}_{kh}(\bar{{\bf{q}}})\|_{L^{\infty}(I)}\|\bar{\mu}_{kh}\|_{L^{1}(\bar{I})}$
-$\leq$ $2\|{\bf w}\|_{L^{2}(\Omega)}\|\bar{{\bf{u}}}(\bar{{\bf{q}}})-({\bf{u}}_{kh}(\bar{{\bf{q}}})\|_{L^{\infty}(I;L^{2}(\Omega))}\|\bar{\mu}_{kh}\|_{L^{1}(\bar{I})}$
+$\leq$ $2\|{\bf w}\|_{L^{2}(\Omega)}\|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{u}}}(\bar{{\bf{q}}})-({\bf{u}}_{kh}(\bar{{\bf{q}}})\|_{L^{\infty}(I;L^{2}(\Omega))}\|\bar{\mu}_{kh}\|_{L^{1}(\bar{I})}$
 $\leq$ $C\ell_{k}^{2}(h^{2}+k)\|\bar{{\bf{q}}}\|_{L^{\infty}(I;L^{2}(\Omega))}\|{\bf w}\|_{L^{2}(\Omega)}\|\bar{\mu}_{kh}\|_{L^{1}(\bar{I})}$
@@
 $+(\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d},{\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}})_{I\times\Omega}+(\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d},\bar{{\bf{u}}}-\bar{{\bf{u}}}_{kh})_{I\times\Omega}$
-$=$ $-\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{kh}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+(\bar{{\bf{u}}}-{\bf{u}}_{d},{\bf{u}}(\bar{{\bf{q}}}_{kh})-\bar{{\bf{u}}}_{kh})_{I\times\Omega}+(\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d},{\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}})_{I\times\Omega}$
+$=$ $-\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{kh}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+(\bar{{\bf{u}}}-{\bf{u}}_{d},{\bf{u}}(\bar{{\bf{q}}}_{kh})-\bar{{\bf{u}}}_{kh})_{I\times\Omega}+(\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d},{\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}})_{I\times\Omega}$
 by using Corollary 4.7, we obtain
-$\alpha$ $\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{kh}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{kh}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}$
+$\alpha$ $\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{kh}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{kh}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$
 $\leq C\ell^{2}_{k}(k+h^{2})+(\bar{{\bf{u}}}-{\bf{u}}_{d},{\bf{u}}(\bar{{\bf{q}}}_{kh})-\bar{{\bf{u}}}_{kh})_{I\times\Omega}+(\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d},{\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}})_{I\times\Omega}$
-$\leq C\ell^{2}_{k}(k+h^{2})+\|\bar{{\bf{u}}}-{\bf{u}}_{d}\|_{L^{2}(I;L^{2}(\Omega))}\|{\bf{u}}(\bar{{\bf{q}}}_{kh})-\bar{{\bf{u}}}_{kh}(\bar{{\bf{q}}}_{kh})\|_{L^{2}(I;L^{2}(\Omega))}$
-$\quad+\|\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d}\|_{L^{2}(I;L^{2}(\Omega))}\|{\bf{u}}_{kh}(\bar{{\bf{q}}})-\bar{{\bf{u}}}(\bar{{\bf{q}}})\|_{L^{2}(I;L^{2}(\Omega))}$
-$\leq C\ell^{2}_{k}(k+h^{2})\left[1+\left(\|\bar{{\bf{u}}}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}+\|{\bf{u}}_{d}\|_{L^{2}(I;L^{2}(\Omega))}\right)\left(\|\bar{{\bf{q}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{{\bf{q}}}\|_{L^{2}(I;L^{2}(\Omega))}\right)\right]$
+$\leq C\ell^{2}_{k}(k+h^{2})+\|\bar{{\bf{u}}}-{\bf{u}}_{d}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\|{\bf{u}}(\bar{{\bf{q}}}_{kh})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{u}}_{kh}}(\bar{{\bf{q}}}_{kh})\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$
+$\quad+\|\bar{{\bf{u}}}_{kh}-{\bf{u}}_{d}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\|{\bf{u}}_{kh}(\bar{{\bf{q}}})-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{u}}}(\bar{{\bf{q}}})\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$
+$\leq C\ell^{2}_{k}(k+h^{2})\left[1+\left(\|\bar{{\bf{u}}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\|\bar{{\bf{u}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\|{\bf{u}}_{d}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\right)\left(\|\bar{{\bf{q}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\|\bar{{\bf{q}}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\right)\right]$
 $\leq C\ell^{2}_{k}(k+h^{2}),$
-where in the last step, we used the boundedness of $\|\bar{{\bf{u}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}$, $\|\bar{{\bf{u}}}\|_{L^{2}(I;L^{2}(\Omega))}$, $\|\bar{{\bf{q}}}\|_{L^{2}(I;L^{2}(\Omega))}$, and $\|\bar{{\bf{q}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}$ from Theorem 3.8 and Lemma 5.7.
+where in the last step, we used the boundedness of $\|\bar{{\bf{u}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$, $\|\bar{{\bf{u}}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$, $\|\bar{{\bf{q}}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$, and $\|\bar{{\bf{q}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$ from Theorem 3.8 and Lemma 5.7.
 ∎
@@
 We introduce the projection $\pi_{d}\colon L^{2}(I;L^{2}(\Omega)^{d})\to\mathbf{Q}_{0}$, defined by
-$\left({\bf{q}}-\pi_{d}{\bf{q}},\mathbf{r}\right)_{I\times\Omega}=0\qquad\forall\mathbf{r}\in\mathbf{Q}_{0}.$ (6.2)
+$\left({\bf{q}}-\pi_{d}{\bf{q}},\mathbf{r}\right)_{I\times\Omega}=0\qquad\forall\mathbf{r}\in\mathbf{Q}_{0},$ (6.2)
+which by definition is stable in ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}$, i.e., satisfies
+$\|\pi_{d}{\bf{q}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq\|{\bf{q}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}.$ (6.3)
 Note that this projection satisfies the explicit formula
@@
 Further, we have $\pi_{d}(\mathbf{Q}_{ad})\subset\mathbf{Q}_{0,ad}$. We can now formulate the fully discrete optimal control problem, which reads
-Minimize $\text{Minimize }\;J({\bf{q}}_{\sigma},{\bf{u}}_{\sigma})=\frac{1}{2}\lVert{\bf{u}}_{\sigma}-{\bf{u}}_{d}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\frac{\alpha}{2}\lVert{\bf{q}}_{\sigma}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}$ (6.4a)
+Minimize $\text{Minimize }\;J({\bf{q}}_{\sigma},{\bf{u}}_{\sigma})=\frac{1}{2}\lVert{\bf{u}}_{\sigma}-{\bf{u}}_{d}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\frac{\alpha}{2}\lVert{\bf{q}}_{\sigma}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$ (6.5a)
 subject to $({\bf{q}}_{\sigma},{\bf{u}}_{\sigma})\in\mathbf{Q}_{0,ad}\times X^{0}_{k}(\mathbf{V}_{h})$, satisfying
@@
 As the discrete solution operator $S_{kh}$ is linear and continuous from $L^{2}(I;L^{2}(\Omega)^{d})\to L^{\infty}(I;L^{2}(\Omega)^{d})$, we have
-$G_{\bf{w}}({\bf{u}}_{kh}(\pi_{d}\tilde{\bf{q}}))=G_{\bf{w}}({\bf{u}}_{kh}(\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}))+G_{\bf{w}}({\bf{u}}_{kh}(\tilde{\bf{q}}))<C\|\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}\|_{L^{2}(I;L^{2}(\Omega))}+\beta-\delta.$
+$G_{\bf{w}}({\bf{u}}_{kh}(\pi_{d}\tilde{\bf{q}}))=G_{\bf{w}}({\bf{u}}_{kh}(\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}))+G_{\bf{w}}({\bf{u}}_{kh}(\tilde{\bf{q}}))<C\|\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\beta-\delta.$
 As 3 guarantees $k\to 0$ as $h\to 0$, it holds $\|\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}\|_{L^{2}(I;L^{2}(\Omega))}\to 0$ for $h\to 0$.
-This implies that for $h$ small enough, we have $C\|\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}\|_{L^{2}(I;L^{2}(\Omega))}\leq\delta$ and as a consequence $G_{\bf{w}}({\bf{u}}_{kh}(\pi_{d}\tilde{\bf{q}}))<\beta$.
+This implies that for $h$ small enough, we have $C\|\pi_{d}\tilde{\bf{q}}-\tilde{\bf{q}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq\delta$ and as a consequence $G_{\bf{w}}({\bf{u}}_{kh}(\pi_{d}\tilde{\bf{q}}))<\beta$.
 ∎
@@
 Then there exists a constant $C>0$ independent of $k$ and $h$, such that the fully discrete optimal control $\bar{{\bf{q}}}_{\sigma}\in\mathbf{Q}_{0,ad}$, solving (6.5a)-(6.5c), together with its corresponding state $\bar{{\bf{u}}}_{\sigma}\in X^{0}_{k}(\mathbf{V}_{h})$ and corresponding multiplier $\bar{\mu}_{\sigma}\in L^{1}(\bar{I})$ satisfies the bound
-$\|\bar{{\bf{q}}}_{\sigma}\|_{L^{\infty}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{\mu}_{\sigma}\|_{(C(\bar{I}))^{*}}\leq C.$
+$\|\bar{{\bf{q}}}_{\sigma}\|_{L^{\infty}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{\sigma}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\|\bar{\mu}_{\sigma}\|_{(C(\bar{I}))^{*}}\leq C.$
 Proof.
 By Lemma 6.1 under the given assumptions, the fully discrete control $\pi_{d}\tilde{\bf{q}}$ is feasible, and thus it holds
-$J(\bar{{\bf{q}}}_{\sigma},\bar{{\bf{u}}}_{\sigma})$ $\leq J(\pi_{d}\tilde{{\bf{q}}},\bar{{\bf{u}}}_{kh}(\pi_{d}\tilde{{\bf{q}}}))$
-$=\frac{1}{2}\lVert\bar{{\bf{u}}}_{kh}(\pi_{d}\tilde{{\bf{q}}})-{\bf{u}}_{d}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\frac{\alpha}{2}\lVert\pi_{d}\tilde{{\bf{q}}}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}$
-$\leq\lVert\bar{{\bf{u}}}_{kh}(\tilde{{\bf{q}}})-{\bf{u}}(\tilde{{\bf{q}}})\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\lVert{\bf{u}}(\tilde{{\bf{q}}})-{\bf{u}}_{d}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\frac{\alpha}{2}\lVert\tilde{{\bf{q}}}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}$
-$\leq C(T,\tilde{{\bf{q}}},{\bf{u}}_{d}).$
-As a result we obtain $\|\bar{{\bf{q}}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{{\bf{u}}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}\leq C.$ The proof of $\|\bar{\mu}_{\sigma}\|_{C(\bar{I})^{*})}\leq C$ and $\|\bar{\bf{q}}_{\sigma}\|_{L^{\infty}(I;L^{2}(\Omega))}\leq C$ is then accomplished by following the same steps as the proof of Lemma 5.7, making use of $G_{\bf{w}}({\bf{u}}_{kh}({\pi_{d}{\bf{p}}}))\leq\beta-\frac{1}{4}\delta$ for $h$ sufficiently small.
+$J(\bar{{\bf{q}}}_{\sigma},\bar{{\bf{u}}}_{\sigma})$ $\leq J(\pi_{d}\tilde{{\bf{q}}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{u}}_{kh}}(\pi_{d}\tilde{{\bf{q}}}))$
+$=\frac{1}{2}\lVert{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{u}}_{kh}}(\pi_{d}\tilde{{\bf{q}}})-{\bf{u}}_{d}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\frac{\alpha}{2}\lVert\pi_{d}\tilde{{\bf{q}}}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$
+$\leq\lVert{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bf{u}}_{kh}}(\pi_{d}\tilde{{\bf{q}}})\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\lVert{\bf{u}}_{d}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\frac{\alpha}{2}\lVert\tilde{{\bf{q}}}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$
+$\leq C(T,\tilde{{\bf{q}}},{\bf{u}}_{d}),$
+where in the last step, we have used the discrete stability result for ${\bf{u}}_{kh}$ from Theorem 4.2 and the stability of $\pi_{d}$ from (6.3).
+As a result we obtain $\|\bar{{\bf{q}}}_{\sigma}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\|\bar{{\bf{u}}}_{\sigma}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq C.$ The proof of $\|\bar{\mu}_{\sigma}\|_{C(\bar{I})^{*})}\leq C$ and $\|\bar{\bf{q}}_{\sigma}\|_{L^{\infty}(I;L^{2}(\Omega))}\leq C$ is then accomplished by following the same steps as the proof of Lemma 5.7, making use of $G_{\bf{w}}({\bf{u}}_{kh}({\pi_{d}{\bf{p}}}))\leq\beta-\frac{1}{4}\delta$ for $h$ sufficiently small.
 Theorem 6.5.
@@
 Then there exists a constant $C>0$ such that it holds
-$\alpha\lVert\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}_{\sigma}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\lVert\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{\sigma}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}\leq Ch^{2}.$
+$\alpha\lVert\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}_{\sigma}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\lVert\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{\sigma}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq Ch^{2}.$
 Proof.
@@
 Adding these two inequalities, we obtain
-$\alpha\|\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}_{\sigma}\|^{2}_{L^{2}(I;L^{2}(\Omega))}\leq(\bar{{\bf{z}}}_{kh}-\bar{{\bf{z}}}_{\sigma},\bar{{\bf{q}}}_{\sigma}-\bar{{\bf{q}}}_{kh})_{I\times\Omega}+(\alpha\bar{{\bf{q}}}_{\sigma}+\bar{{\bf{z}}}_{\sigma},\pi_{d}\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}_{kh})_{I\times\Omega}:=I_{1}+I_{2}.$ (6.9)
+$\alpha\|\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}_{\sigma}\|^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq(\bar{{\bf{z}}}_{kh}-\bar{{\bf{z}}}_{\sigma},\bar{{\bf{q}}}_{\sigma}-\bar{{\bf{q}}}_{kh})_{I\times\Omega}+(\alpha\bar{{\bf{q}}}_{\sigma}+\bar{{\bf{z}}}_{\sigma},\pi_{d}\bar{{\bf{q}}}_{kh}-\bar{{\bf{q}}}_{kh})_{I\times\Omega}:=I_{1}+I_{2}.$ (6.10)
 We estimate the two terms separately.
@@
 $=(\bar{{\bf{u}}}_{\sigma}-\bar{{\bf{u}}}_{kh},\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{\sigma})_{I\times\Omega}+\langle G_{\bf w}(\bar{{\bf{u}}}_{\sigma})-G_{\bf w}(\bar{{\bf{u}}}_{kh}),\bar{\mu}_{kh}\rangle-\langle G_{\bf w}(\bar{{\bf{u}}}_{\sigma})-G_{\bf w}(\bar{{\bf{u}}}_{kh}),\bar{\mu}_{\sigma}\rangle$
-$\leq-\lVert\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{\sigma}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}+\langle\beta-G_{\bf w}(\bar{{\bf{u}}}_{kh}),\bar{\mu}_{kh}\rangle+\langle\beta-G_{\bf w}(\bar{{\bf{u}}}_{\sigma}),\bar{\mu}_{\sigma}\rangle$
-$=-\lVert\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{\sigma}\rVert^{2}_{L^{2}(I;L^{2}(\Omega))}.$
+$\leq-\lVert\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{\sigma}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\langle\beta-G_{\bf w}(\bar{{\bf{u}}}_{kh}),\bar{\mu}_{kh}\rangle+\langle\beta-G_{\bf w}(\bar{{\bf{u}}}_{\sigma}),\bar{\mu}_{\sigma}\rangle$
+$=-\lVert\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{\sigma}\rVert^{2}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}.$
 By the Cauchy-Schwarz inequality and properties of the $L^{2}$-projection
@@
 $=\left(\bar{\bf{z}}_{\sigma}-\pi_{d}\bar{\bf{z}}_{\sigma},\pi_{d}\bar{\bf{q}}_{kh}-\bar{\bf{q}}_{kh}\right)_{I\times\Omega}$
-$\leq Ch^{2}\|\nabla\bar{\bf{z}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}\|\nabla\bar{\bf{q}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}.$
-Using that $\|{\rm NewA}\bar{{\bf{q}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}\leq\alpha^{-1}\|{\rm NewA}\bar{{\bf{z}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}$, the stability of $P_{ad}$ and the stability of solutions to the fully discrete dual problem, pointed out in Remark 4.3, gives
-$I_{2}\leq Ch^{2}\left(\|\bar{{\bf{u}}}_{\sigma}-\bar{{\bf{u}}}_{d}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{\mu}_{\sigma}\|_{L^{1}(I)}\|{\bf{w}}\|_{L^{2}(\Omega)}\right)\left(\|\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{d}\|_{L^{2}(I;L^{2}(\Omega))}+\|\bar{\mu}_{kh}\|_{L^{1}(I)}\|{\bf{w}}\|_{L^{2}(\Omega)}\right).$
-Now the boundedness of $\|\bar{{\bf{u}}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}$, $\|\bar{{\bf{u}}}_{kh}\|_{L^{2}(I;L^{2}(\Omega))}$, $\|\bar{\mu}_{\sigma}\|_{L^{1}(I)}$, and $\|\bar{\mu}_{kh}\|_{L^{1}(I)}$ from Lemmas 5.7 and 6.4 finish the proof.
+$\leq Ch^{2}\|\nabla\bar{\bf{z}}_{\sigma}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\|\nabla\bar{\bf{q}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}.$
+Using that $\|{\rm NewA}\bar{{\bf{q}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq\alpha^{-1}\|{\rm NewA}\bar{{\bf{z}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$, which holds due to the projection formula (5.5d) and the stability of $P_{ad}$ in $H^{1}(\Omega)$, see [27, Theorem A.1], [2, Theorem 5.8.2], and the stability of solutions to the fully discrete dual problem, pointed out in Remark 4.3, gives
+$I_{2}\leq Ch^{2}\left(\|\bar{{\bf{u}}}_{\sigma}-\bar{{\bf{u}}}_{d}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\|\bar{\mu}_{\sigma}\|_{L^{1}(I)}\|{\bf{w}}\|_{L^{2}(\Omega)}\right)\left(\|\bar{{\bf{u}}}_{kh}-\bar{{\bf{u}}}_{d}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\|\bar{\mu}_{kh}\|_{L^{1}(I)}\|{\bf{w}}\|_{L^{2}(\Omega)}\right).$
+Now the boundedness of $\|\bar{{\bf{u}}}_{\sigma}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$, $\|\bar{{\bf{u}}}_{kh}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$, $\|\bar{\mu}_{\sigma}\|_{L^{1}(I)}$, and $\|\bar{\mu}_{kh}\|_{L^{1}(I)}$ from Lemmas 5.7 and 6.4 finish the proof.
 ∎
@@
 Let $\bar{\bf{q}}\in\mathbf{Q}_{ad}$ and $\bar{\bf{q}}_{\sigma}\in\mathbf{Q}_{0,ad}$ be the solutions to the continuous and fully discrete optimal control problems (1.1a)-(1.1d) and (6.5a)-(6.5c).
-$\sqrt{\alpha}\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{\sigma}\rVert_{L^{2}(I;L^{2}(\Omega))}+\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{\sigma}\rVert_{L^{2}(I;L^{2}(\Omega))}\leq C\,\ell_{k}(k^{\frac{1}{2}}+h),\quad\ell_{k}=\ln{\frac{T}{k}}.$
+$\sqrt{\alpha}\lVert\bar{{\bf{q}}}-\bar{{\bf{q}}}_{\sigma}\rVert_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}+\lVert\bar{{\bf{u}}}-\bar{{\bf{u}}}_{\sigma}\rVert_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}\leq C\,\ell_{k}(k^{\frac{1}{2}}+h),\quad\ell_{k}=\ln{\frac{T}{k}}.$
 Remark 6.7.
@@
 8 Numerical results
-In the following, we present two numerical examples, studying the orders of convergence for the optimal control problem.
+In the following, we present three numerical examples, studying the orders of convergence for the optimal control problem.
 We first present an example for smooth data.
@@
 In the second example, we study an example where less regularity of the data is available, leading to a reduced order of convergence.
+Lastly, we consider an example with simultaneous control and state constraints.
+Throughout this section, we will analyze the errors of the full discretization $\bar{\bf{q}}-\bar{\bf{q}}_{\sigma}$.
+For details regarding the implementation of a variational discretization in the presence of control constraints, we refer to [21], [22, Chapter 3.2.5] and the references therein.
 8.1 Solution Algorithm
@@
 We discretize this problem with a uniform triangulation of $\Omega$ and a uniform partition of $I$.
-To get more insight into the observed orders of convergence, for sequences of discretization levels $\{h_{l}\}$,$\{k_{l}\}$ we report the empirical orders of convergence determined by
-$(\mathrm{EOC}_{h})_{l}$ $:=\dfrac{\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{k_{l}h_{l}}\|_{L^{2}(I;L^{2}(\Omega))})-\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{k_{l-1}h_{l-1}}\|_{L^{2}(I;L^{2}(\Omega))})}{\log(h_{l})-\log(h_{l-1})}$
-$(\mathrm{EOC}_{k})_{l}$ $:=\dfrac{\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{k_{l}h_{l}}\|_{L^{2}(I;L^{2}(\Omega))})-\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{k_{l-1}h_{l-1}}\|_{L^{2}(I;L^{2}(\Omega))})}{\log(k_{l})-\log(k_{l-1})}$
-Figure 1(a) displays convergence with respect to the spacial discretization parameter for fixed $k$.
+To get more insight into the observed orders of convergence, for a sequence of discretization levels $\{\sigma_{l}\}=\{(h_{l},k_{l})\}$, we report the empirical orders of convergence determined by
+$(\mathrm{EOC}_{h})_{l}$ $:=\dfrac{\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{l}}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}})-\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{l-1}}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}})}{\log(h_{l})-\log(h_{l-1})}$
+$(\mathrm{EOC}_{k})_{l}$ $:=\dfrac{\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{l}}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}})-\log(\|\bar{\bf{q}}-\bar{\bf{q}}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{l-1}}}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}})}{\log(k_{l})-\log(k_{l-1})}$
+Figure 3 displays convergence with respect to the spacial discretization parameter for fixed $k$.
 The theoretical convergence order of 1 can be observed.
-Figure 1(b) depicts the convergence results for the time discretization parameter for different values of $h$.
+Figure 3 depicts the convergence results for the time discretization parameter for different values of $h$.
 Note that for this comparison, we have choosen discretization levels in time, such that the two boundary of the active set $[1/4,3/4]$ are midpoints of the discrete subintervals, in order to exclude superconvergence effects.
 Due to the structure of the discretization, the number of degrees of freedom, grows rather quickly, which is why very fine discretizations are expensive.
-Figure 1(b) shows, that if the spacial discretization parameter is chosen too large, the stagnation phase sets in rather early.
+Figure 3 shows, that if the spacial discretization parameter is chosen too large, the stagnation phase sets in rather early.
 If one only observes the coarse discretizations, the observed order of convergence is skewed, which led to an estimated order of convergence of 0.85 reported in [8].
@@
 $2^{-6}$ 3.9657 1.02
-((a)) Convergence with respect to $h$ for $k=10^{-3}$.
+Figure 1: Convergence with respect to $h$ for $k=10^{-3}$.
 $h=2^{-6}$ $h=2^{-7}$ $h=2^{-8}$
@@
 $162^{-1}$ 4.0066 0.07 2.0579 0.22 1.1443 0.50
-((b)) Convergence with respect to $k$ for different values of $h$.
-Figure 1: Numerical observation of the error $\|\bar{\bf{q}}-\bar{\bf{q}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}$ for Example 1.
+Figure 2: Convergence with respect to $k$ for different values of $h$.
+Figure 3: Numerical observation of the error $\|\bar{\bf{q}}-\bar{\bf{q}}_{\sigma}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$ for Example 1.
 8.3 Example 2
@@
 Note that due to this evaluation of the errors, we expect a faster convergence than theoretically derived, as on the finest discretization level, the error would equal to 00.
-In Figure 3(a) we can again observe order 1 convergence with respect to $h$ for fixed time discretization.
-In Figure 3(b) we observe that the convergence with respect to $k$ exhibits a rate of about $0.6$, which is much closer to the analytically derived $0.5$ than the rate observed for the smooth example 1.
-((a)) Time function $\varphi(t)$.
-((b)) Weight function ${\bf{w}}$.
-((c)) Satisfaction of the state constraint.
+In Figure 10 we can again observe order 1 convergence with respect to $h$ for fixed time discretization.
+In Figure 10 we observe that the convergence with respect to $k$ exhibits a rate of about $0.6$, which is much closer to the analytically derived $0.5$ than the rate observed for the smooth example 1.
+Figure 4: Time function $\varphi(t)$.
+Figure 5: Weight function ${\bf{w}}$.
+Figure 6: Satisfaction of the state constraint.
 Figure 7: Data for Example 2.
@@
 $2^{-5}$ 42.372 1.18
-((a)) Convergence with respect to $h$ for $k=960^{-1}$.
+Figure 8: Convergence with respect to $h$ for $k=960^{-1}$.
 $k$ err $\mathrm{EOC}_{k}$
@@
 $80^{-1}$ 74.697 0.66
-((b)) Convergence with respect to $k$ for $h=2^{-7}$.
-Figure 3: Numerical observation of the error $\|\bar{\bf{q}}-\bar{\bf{q}}_{\sigma}\|_{L^{2}(I;L^{2}(\Omega))}$ for Example 2.
+Figure 9: Convergence with respect to $k$ for $h=2^{-7}$.
+Figure 10: Numerical observation of the error $\|\bar{\bf{q}}-\bar{\bf{q}}_{\sigma}\|_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{2}(I\times\Omega)}}$ for Example 2.
+8.4 Example 3
+In the previous examples, no control constraints were present.
+To highlight, that the derived error estimates are indeed not influenced by the control constraints, we augment the Example of Section 8.2 by a control constraint $\bar{\bf{q}}\leq 200\cdot\mathds{1}$.
+The remaining choices of $\Omega,{\bf{w}},{\bf{u}}_{d},\alpha$ and $\beta$ are kept the same.
+Due to the presence of the control constraint, an analytic solution $\bar{\bf{q}}$ to the optimal control problem is not known, and we again compare to a fine-grid solution, computed with $k=960^{-1}$ and $h=2^{-7}$.
+The numerically observed orders of convergence in $k$ and $h$, as well as the active set of the control contstraint for a fixed poin in time can be observed in Figure 11.
+It has to be noted, that while the discrete problem has a similar structure to the one without control constraint, and can be solved by the same PDAS algorithm, the resulting saddle point systems can be much larger for large active sets.
+Efficient preconditioners for such problems are needed, to study the performance of this algorithm for finer discretizations.
+$k=960^{-1}$
+$h$ err $\mathrm{EOC}_{h}$
+$2^{-2}$ 64.049 -
+$2^{-3}$ 35.438 0.85
+$2^{-4}$ 16.024 1.14
+$2^{-5}$ 6.8962 1.21
+$h=2^{-7}$
+$k$ err $\mathrm{EOC}_{k}$
+$5^{-1}$ 18.915 -
+$10^{-1}$ 9.1295 1.05
+$20^{-1}$ 4.5816 0.99
+$40^{-1}$ 2.3133 0.98
+$80^{-1}$ 1.1640 0.99
+$160^{-1}$ 0.5827 0.99
+Figure 11: Left: Orders of convergence for an example with control constraints.
+Right: Active sets of the discrete optimal control for $h=2^{-5}$, $k=30^{-1}$ at $t=1/3$.
+(grey: inactive, blue: constraint in $x_{1}$ direction active, red: constraint in $x_{2}$ direction active).